### v2

# Spectral Criteria for the Asymptotics of Local Functionals of Gaussian Fields and Their Application to Nodal Volumes

Abstract
In this paper, we establish a general criterion for the positivity of the variance of the second chaotic component of local functionals of stationary vector-valued Gaussian fields. This criterion is formulated in terms of the spectral properties of the covariance function, without requiring assumptions of integrability or isotropy. It offers a simple and robust framework for analyzing variance asymptotics in such models. We apply this approach to the study of the nodal volume and the number of critical points of a Gaussian field, proving the positivity of the limiting variance under mild conditions on the covariance function. Additionally, we examine the asymptotics of the particular model of Euclidean random waves, deriving the central limit theorem through an analysis of the second and fourth chaotic components, thereby resolving an open question in large dimensions. As a byproduct, we condense and generalize many existing results on the volume of intersections of random waves, bypassing the need for traditional, intricate variance computations. Our findings shed new light on the second-chaos cancellation phenomenon from a spectral perspective and can be extended to any local, possibly singular, functional of Gaussian fields.
00footnotetext: University of Luxembourg00footnotetext: This work was supported by the Luxembourg National Research Fund (Grant: 021/16236290/HDSA)00footnotetext: Email: louis.gass(at)uni.lu
1 Introduction and main results
1.1 A general framework
Let $V$ and $W$ be two finite dimensional vector spaces of dimension $d$ and $n$ respectively. Let $X:V\rightarrow W$ be a vector-valued stationary Gaussian random field with a continuous non-degenerate covariance function $\Omega:V\rightarrow S^{2}(W^{*})$, where this last space denotes the set of symmetric bilinear map on $W$. We denote by $\eta$ the associated Gaussian measure on $W$, and let $f:W\rightarrow\mathbb{R}$ in $L^{2}(\eta)$. Our primal object of interest is the quantity
$Z_{\lambda}(\phi)=\frac{1}{\lambda^{d/2}}\int_{V}\phi\left(\frac{v}{\lambda}\right)f(X(v))\mathrm{d}v,$
for a test function $\phi$ such that $\|\phi\|_{2}=1$. We say that $Z_{\lambda}$ satisfies a Central Limit Theorem (CLT) if the following convergence in distribution holds
$\frac{Z_{\lambda}(\phi)-\mathbb{E}[Z_{\lambda}(\phi)]}{\sqrt{\operatorname{Var}(Z_{\lambda}(\phi))}}\underset{\lambda\rightarrow+\infty}{\longrightarrow}\mathcal{N}(0,1)$
A proof of the CLT is usually split into two very distinct problems: the asymptotic normality and the positivity of the limiting variance. The asymptotic normality is well-understood and was first proved in [Cuz76] in the case $d=n=1$ and $\Omega\in L^{1}(\mathbb{R}^{d})$ by approximation using $m$-dependent processes. Since then, it has been refined with the introduction of the Wiener chaos expansion, which has become an indispensable tool for proving asymptotics related to $Z_{\lambda}(\phi)$. In particular, the celebrated Breuer–Major theorem [Bre83] states that asymptotic normality holds if the Hermite rank of $f$ is $q$ and $\Omega\in L^{q}(V)$. Modern proofs of this theorem now include quantitative CLT and functional convergence [Nou11, Nou20, Nua20] and make use of more advanced techniques, such as Malliavin calculus and Stein’s method. The literature on the topic is extensive.
In the case $n=1$, the positivity of the limiting variance can be established using simple Fourier arguments under minimal conditions, such as assuming that $f$ is not odd or that the spectral measure has a density with respect to the origin, see [Cuz76]. For $n>1$, the situation is more intricate, and the positivity of the limiting variance cannot be inferred from straightforward observations on $f$ and $\Omega$. The first goal of this paper is to establish general conditions on $\Omega$ and $f$ to ensure that the limiting variance of $Z_{\lambda}(\phi)$ is positive, i.e., that $Z_{\lambda}(\phi)$ does not vanish at infinity.
The heuristic of the approach for obtaining a lower bound for the limiting variance is the following. Since the process $f\circ X:V\rightarrow\mathbb{R}$ is stationary, its covariance function is a positive semi-definite function, meaning that its Fourier transform $\rho$ is a finite measure on $V^{*}$. This fact can in turn be exploited to obtain lower bound for the variance in (1.1). Indeed, in the simple case where $\rho$ has a density near the origin w.r.t the Lebesgue measure then
$\lim_{\lambda\rightarrow+\infty}\operatorname{Var}(Z_{\lambda}(\phi))=\rho(0),$
thus the limiting variance is positive if and only if $\rho(0)$ is positive. In a more general framework, the variance can be understood from local properties of the spectral measure $\rho$ near the origin. To explicitly compute the measure $\rho$ in terms of the spectral measure $\mu$ associated to $X$, the most direct path is to expand $f$ into its orthogonal Hermite decomposition on the space $L^{2}(\eta)$. The variance will then be expressed as sum of powers of the covariance function $\Omega$. Passing to Fourier transform, this yields an expression of $\rho$ in terms of a sum of non-negative terms, given by iterated convolutions of the spectral measure $\mu$ and Hermite expansion of $f$. Additional comments on the advantage of this approach will be made after Theorem 1.4.
[…]
$f=\sum_{q=0}^{+\infty}\pi_{q}(f)=\sum_{q=0}^{+\infty}\frac{1}{q!}\langle f_{q},H^{q}\rangle.$
We refer to Section 2 for more details.
Since in the following, we use Fourier transform arguments, but we want to keep the notations to general vector space notations, we introduce $W_{\mathbb{C}}=W+iW$, the complexification of $W$. One must simply understand $W$ as $\mathbb{R}^{n}$ and $W_{\mathbb{C}}$ as $\mathbb{C}^{n}$. By Bochner theorem, the Fourier transform $\mu$ of the covariance function $\Omega$ is an operator-valued measure on $V^{*}$, taking values in the space of positive Hermitian form on $W_{\mathbb{C}}^{*}$. Its $q$-fold convolution, denoted by $\mu^{*q}$, can be seen as measure on $V$, taking values in the space of positive Hermitian forms on the space of symmetric $q$-linear forms on $W_{\mathbb{C}}$. These facts are explained in more detail in Section 3.
In the following, we fix a function $\phi\in L^{1}\cap L^{2}(V)$, with $\|\phi\|_{2}=1$, and we define $\gamma=|\hat{\phi}|^{2}$ on $V^{*}$, where the $\hat{}$ notation stand for the Fourier transform. Let $\gamma_{\lambda}:v\mapsto\lambda^{d}\gamma(\lambda v)$, which weakly asymptotically converges to the Dirac delta function at 00, since by Plancherel isometry, $\|\gamma\|_{1}=1$. We define the least radially decreasing majorant and the greatest radially decreasing minorant of $\gamma$ by the formulas
$\gamma^{+}:x\rightarrow\sup_{\|y\|\geq\|x\|}\gamma(y)\quad\text{and}\quad\gamma^{-}:x\rightarrow\inf_{\|y\|\leq\|x\|}\gamma(y).$
We say that $\phi$ is a test function if $\gamma^{+}\in L^{1}(V^{*})$. By property of the Fourier transform, if the function $\phi$ is absolutely continuous with bounded variation, or if it is the Fourier transform of a ball or more generally of a smooth compact hypersurface with positive Gauss curvature, this hypothesis is satisfied, see [Her62] and the heuristic explanation before Lemma 4.6. Note also that $\gamma$ is continuous and $\gamma(0)=1$, so that $\gamma^{-}$ is nonzero.
1.1.1 A general spectral expression for the variance
We define the $q$-th chaotic projection of $Z_{\lambda}(\phi)$ by
$Z_{\lambda}^{(q)}(\phi)=\frac{1}{\lambda^{d/2}}\int_{V}\phi\left(\frac{v}{\lambda}\right)\pi_{q}(f)(X(v))\mathrm{d}v,$
{mdframed}
[backgroundcolor=white]
Theorem 1.1.
One has
$\operatorname{Var}(Z_{\lambda}(\phi))=\sum_{q=1}^{+\infty}\operatorname{Var}(Z_{\lambda}^{(q)}(\phi)).$
and
$\operatorname{Var}(Z_{\lambda}^{(q)}(\phi))=\frac{1}{q!}\int_{V^{*}}\gamma_{\lambda}(\xi)\,\mathrm{d}\mu^{*q}(f_{q})(\xi).$
Since $\mu^{*q}$ is a positive Hermitian measure, the measure $\mu^{*q}(f_{q})$ is non-negative. The expression of the variance in 1.1 then has the advantage, compared to the "non-spectral" expression for the variance of the $q$-th chaos, of being written as the integrand of a non-negative measure, which is our first starting point to give lower bound for the limiting variance. As $\lambda$ goes to infinity, this expression formally converges to the Lebesgue density at 00 of the measure $\mu^{*q}(f_{q})$. To precise this heuristic, we define for a real bounded measure $\nu$ on $V^{*}$ the quantities
$D^{+}\nu=\limsup_{R\rightarrow 0}\frac{\nu(B(0,R))}{\operatorname{Vol}(B(0,R))}\quad\text{and}\quad D^{-}\nu=\liminf_{R\rightarrow 0}\frac{\nu(B(0,R))}{\operatorname{Vol}(B(0,R))}.$
If $D^{-}\nu=D^{+}\nu$, we call this quantity $D\nu$. The quantity $D\nu$, when it exists, can be thought as a weak version of the density at 00 of the measure $\nu$ w.r.t the Lebesgue measure on $V^{*}$. {mdframed}[backgroundcolor=white]
Corollary 1.2.
$\liminf_{\lambda\rightarrow+\infty}\operatorname{Var}(Z_{\lambda}^{(q)}(\phi))\geq\frac{1}{q!}\|\gamma^{-}\|_{1}D^{-}(\mu^{*q}(f_{q})),$
and
$\limsup_{\lambda\rightarrow+\infty}\operatorname{Var}(Z_{\lambda}^{(q)}(\phi))\leq\frac{1}{q!}\|\gamma^{+}\|_{1}D^{+}(\mu^{*q}(f_{q})).$
This corollary, which is a direct corollary of Theorem 1.1 and Lemma 3.8, puts forward the interaction between the iterated convolutions of the spectral measure, their regularity near the origin, and asymptotic results of local functionals of the Gaussian field. In the case $n=1$, the $q$-fold convolution of $\mu$ can be seen as the distribution of a random walk with transition probability distribution $\mu$ after $q$ steps. The properties of $\mu^{*q}$ near the origin are linked to the probability of return around 00 of the random walk. This analogy has been successfully used in [Lac22] to obtain variance asymptotics related to the volume of random excursions sets, and in [Gro24] to study the covariance structure of the Berry random wave model.
If $D\nu$ exists and is finite, one can conclude for general a measure $\mu$ that the variance converges if the function $\gamma$ is radially decreasing. If one assumes more hypothesis on $\mu^{*q}$, for instance that it has a bounded continuous density, then the convergence follows from more classical results concerning convolution. This is for instance the case if the covariance function $\Omega$ is in $L^{q}$, which is the classical setting of the Breuer-Major theorem, and one can drop the assumption that $\gamma^{+}$ is integrable. The following theorem gives alternative expressions in the case where the spectral measure $\mu$ has a density in $L^{p}$, where $p$ is the conjugate exponent of $q$. {mdframed}[backgroundcolor=white]
Corollary 1.3.
Let $p$ be the conjugate exponent to $q$. If $\mu$ has a density $\Sigma\in L^{p}(V^{*})$ (in $C_{0}(V^{*})$ if $q=1$) w.r.t the Lebesgue measure, then
$\lim_{\lambda\rightarrow+\infty}\operatorname{Var}(Z_{\lambda}^{(q)}(\phi))=\frac{1}{q!}\int_{\xi_{1}+\ldots+\xi_{q}=0}\bigotimes_{k=1}^{q}\Sigma(\xi_{k})(f_{q})\mathrm{d}\xi.$
For $q=1$,
$\lim_{\lambda\rightarrow+\infty}\operatorname{Var}(Z_{\lambda}^{(1)}(\phi))=\Sigma(0)(f_{1}).$
For $q=2$, in matrix form,
$\lim_{\lambda\rightarrow+\infty}\operatorname{Var}(Z_{\lambda}^{(2)}(\phi))=\frac{1}{2}\int_{V^{*}}\operatorname{Tr}(\Sigma(\xi)f_{2}\Sigma(\xi)f_{2})\mathrm{d}\xi.$
1.1.2 A general cancellation theorem for the variance
If $n=1$ then $f_{q}$ is a real number and $\mu$ is a bounded measure on $V^{*}$, so that $\mu^{*q}(f_{q})=f_{q}^{2}\mu^{*q}$. At a fixed chaos, if $f_{q}$ is nonzero, the variance asymptotics is solely contained in the spectral measure $\mu$. This observation is substantially wrong as soon as the dimension $n$ is greater than one. It can happen that the variance asymptotics greatly differs for two different non zero $q$-linear maps $f_{q}$ and $g_{q}$. This general fact is generally described as a cancellation phenomenon. This phenomenon is remarkable in the sense that a cancellation phenomenon can happen only for a class of $q$-linear map of codimension at least one, i.e. almost never.
Several examples of cancellation phenomenon are issued from the context of random waves, see Section 1.3, but let us give a simple one here. Let $Y(t)$ be a random Gaussian process with covariance function $r(t)=\sin(t)/t$. We define the Gaussian process $X:\mathbb{R}\rightarrow\mathbb{R}^{2}$ by $X(t)=(Y(t),Y(t+\pi))$, with covariance function $\Omega$. Let $f,g:\mathbb{R}^{2}\rightarrow\mathbb{R}$ be defined by
$f(x,y)=(x^{2}-1)-(y^{2}-1)\quad\text{and}\quad g(x,y)=(x^{2}-1)+(y^{2}-1).$
Both $f$ and $g$ belongs to the second Wiener chaos, and a direct computation shows that
$\lim_{\lambda\rightarrow+\infty}\operatorname{Var}\left(\frac{1}{\sqrt{\lambda}}\int_{0}^{\lambda}f(X(t))\mathrm{d}t\right)=0\quad\text{but}\quad\lim_{\lambda\rightarrow+\infty}\operatorname{Var}\left(\frac{1}{\sqrt{\lambda}}\int_{0}^{\lambda}g(X(t))\mathrm{d}t\right)=4\|r\|_{2}^{2}>0,$
so that the random process $Y$ is subject to cancellation.
For a Hermitian form $f$ on a complex vector space $W$, we define its isotropic cone
$C(f)=\left\{w\in W\leavevmode\nobreak\ \middle|\leavevmode\nobreak\ f(w)=0\right\}.$
The integrand in the limiting variance Corollary 1.3 is non-negative. If there is cancellation, then the integrand is null for almost all coordinates $\xi_{1},\ldots,\xi_{q}$ with null barycenter, which is a very restrictive condition. For instance, in the case $q=1$, the limiting variance cancels if and only if $f\in\operatorname{Ker}(\Sigma(0))$. In the case $q=2$, one can see (see Lemma 3.4 for a proof), that the limiting variance cancels if and only if a.s. $\operatorname{Im}(\Sigma(\xi))\subset C(f_{2})$. More generally, for $q$ even, one can exploit the fact that $\mu^{*2q}=(\mu^{*q})^{*2}$, so that formally, the Lebesgue density at zero of $\mu^{*2q}$ can be thought as the $L^{2}$ norm of $\mu^{*q}$. We define one more concept before stating the last theorem of this section.
Given any measure $\mu$ taking values in the space of positive Hermitian forms on $W$, one can find a real measure $\nu$ on $V^{*}$ such that the measure $\mu$ has a density $\Sigma$ w.r.t to $\nu$, where $\Sigma$ is defined $\nu$-a.s. as a mapping from $V$ to the space of positive Hermitian forms on $w$ (a canonical choice for $\nu$ is the trace of $\mu$). The writing $\mathrm{d}\mu=\Sigma\mathrm{d}\nu$ will be called a representation of $\mu$. One can thus define, the image of $\mu$, denoted by $\operatorname{Im}(\mu)$, as the smallest closed subset $A\subset W^{*}$ such that
$\nu\left(\left\{x\in V\leavevmode\nobreak\ \middle|\leavevmode\nobreak\ \operatorname{Im}(\Sigma(x))\not\subset A\right\}\right)=0.$
This notion coincides with the essential range of the measurable function $(x,w)\mapsto\Sigma(x)(w,.)$, and is independent on a choice of representation of $\mu$, see Lemma 3.7. For $q\geq 0$, the $2q$-th chaotic projection $f_{2q}$ can be seen as an Hermitian form on the dual of symmetric $q$-linear form on $W$, so that the spaces $\operatorname{Im}(\mu^{*q})$ and $C(f_{2q})$ are both subsets of the dual of the space of symmetric $q$-linear map on $W_{\mathbb{C}}$. {mdframed}[backgroundcolor=white]
Theorem 1.4.
  • If $\operatorname{Im}(\mu^{*q})\not\subset C(f_{2q})$ then
    $\liminf_{\lambda\rightarrow+\infty}\operatorname{Var}(Z_{\lambda}^{(2q)}(\phi))>0.$
  • If $\operatorname{Im}(\mu^{*q})\subset C(f_{2q})$, and either
    • $\Omega\in L^{2q}(V)$
    • $\mu^{*q}$ has a $\mathcal{C}^{1}$ density w.r.t the Lebesgue measure of a smooth compact hypersurface,
    then
    $\lim_{\lambda\rightarrow+\infty}\operatorname{Var}(Z_{\lambda}^{(2q)}(\phi))=0.$
Let $\mu^{*q}=\Sigma_{q}\mathrm{d}\nu$ be a representation of $\mu^{*q}$. From Lemma 3.3, one has the equivalence
$\operatorname{Im}(\mu^{*q})\subset C(f_{2q})\quad\Longleftrightarrow\quad\nu-\text{a.s.},\;\;\Sigma_{q}^{\otimes 2}(f_{2q},f_{2q})=0,$
and we will use both forms in applications of Theorem 1.4. In the previous example (1.1.2), the image of the Fourier transform of the covariance matrix $\Omega$ is given by
$\Sigma(\xi)=\begin{pmatrix}\hat{r}(\xi)&\hat{r}(\xi)e^{i\pi\xi}\\ \hat{r}(\xi)e^{-i\pi\xi}&\hat{r}(\xi)\end{pmatrix}\quad\text{and}\quad\operatorname{Im}(\Sigma)=\left\{\alpha\begin{pmatrix}1\\ e^{-i\pi\xi}\end{pmatrix}\leavevmode\nobreak\ \middle|\leavevmode\nobreak\ \alpha\in\mathbb{C},\xi\in\mathbb{R}\right\},$
which included in the isotropic cone of $f$ equal to $C(f)=\left\{(z_{1},z_{2})\in\mathbb{C}^{2}\leavevmode\nobreak\ \middle|\leavevmode\nobreak\ |z_{1}|=|z_{2}|\right\}$, but not in that of $g$, which is reduced to $\{0\}$.
Let us give a short description, in parallel with related works, of the possible behavior of the variance as $\lambda$ goes to infinity of the total variance of $Z_{\lambda}(\phi)$. In the following, we set $q$ to be the Hermite rank of $f$, that is the smallest non-zero index such that $f_{q}$ is nonzero.
– If $\Omega\in L^{q}(V)$ :
  • $\bullet$
    If $q$ is even and $\operatorname{Im}(\mu^{*q/2})\not\subset C(f_{q})$, then Theorem 1.4 is an equivalence and the limiting variance is positive. By the standard contraction principle and Arcones inequality, the variance converges and the CLT holds for $Z_{\lambda}(\phi)$.
  • $\bullet$
    If $q$ is even and $\operatorname{Im}(\mu^{*q/2})\subset C(f_{q})$ then the limiting variance cancels. One must look at the next non-zero chaotic component, if it exists, and prove the positivity of the limit variance, either by the same theorem applied with a greater $q$ or by other mean.
  • $\bullet$
    If $q$ is odd, one must use the general corollary 1.2 or the explicit representation 1.3 if in addition $\mu\in L^{p}$ (this not automatic if $q\geq 3$), to prove a positive lower bound for the variance, from which would follow a CLT.
– If $\Omega\notin L^{q}(V)$ :
In that case, the situation is delicate, and several behaviors may occur.
  • $\bullet$
    In the case $W=\mathbb{R}$, if $\Omega$ is oscillating, meaning that it is not square integrable but the spectral measure has a singularity near the origin which is sufficiently integrable, then it has been proved in [Mai24] that the $q$-th chaos to dominates and the model tends to exhibit a CLT with a variance of higher order than in the case $\Omega\in L^{q}$.
  • $\bullet$
    On the contrary, if $\Omega$ slowly varying, implying in particular that the spectral measure has a singularity at 00 that explodes with a sufficiently bad rate, then the $q$-th chaos again tends to dominate, with a variance growing at a higher order than the case $\Omega\in L^{q}$, but this time exhibits an explicit non-CLT behavior, see for instance [Dob79], and the functional limit is not a Brownian motion. Note that a unifying spectral condition of these two last two points is still missing in the literature, and it is not clear whether the distinction between them is solely contained in the singularity near zero of the spectral measure. One direction toward such result, would be, in light of the fourth moment Theorem [Nua05] and the contraction principle, to write the contractions in term of the spectral measure and observe their behaviors in the limit. These phenomenon are likely to have a multidimensional counterpart.
  • $\bullet$
    If $q$ is even and $\operatorname{Im}(\mu^{*q/2})\subset C(f_{q})$, there is a competition between the cancellation phenomenon and the higher growth of the variance explained in the first two points. From Theorem 1.4, if the inclusion is sufficiently "smooth" then the $q$-th chaos cancels. It might not be the case for more irregular measures, where one can only rely on the general Corollary 1.3. The random waves model, whose spectral measure is supported on the sphere, is a typical model susceptible to fall is this scenario. To obtain the variance asymptotics if the cancellation occurs, one must look at the next non-zero chaotic component and see if it falls in one of the previous aforementioned situations. In any case, the variance asymptotic will be of lower order than the one expected from the heuristic of the case $W=\mathbb{R}$. We refer to Section 1.4 for more details.
1.2 Nodal volume of Gaussian fields
Let $V$ and $U$ be two finite dimensional Euclidean spaces of dimension $d,k$ respectively, with $k\leq d$. Let $Y:V\rightarrow U$ be a non-degenerate stationary Gaussian random field with covariance function $r:V\rightarrow U\otimes U$, and spectral measure $\psi$. We assume that
  • $Y$ is a $\mathcal{C}^{2}$ random field
  • For $u,v\in V$ distinct, the vector $(Y(u),Y(v))$ is non-degenerate
  • $r$ decreases to zero at infinity
It follows from this set of assumption that for $t\in U$, the level set $\{Y=u\}$ is generically a smooth submanifold of codimension $k$ in $V$, to which we associate its random nodal measure $\mathrm{d}Z^{u}$. If $u=0$ we will denote it by $\mathrm{d}Z$. We are interested, in the quantity
$Z_{\lambda}^{u}(\phi)=\frac{1}{\lambda^{d/2}}\int_{V}\phi\left(\frac{v}{\lambda}\right)\mathrm{d}Z^{u}(v).$
The study of the number of zeros of a random function goes back to Kac and Rice [Kac43, Ric45], and has since been intensively studied in the past decades. By means of the Kac–Rice formula, the quantity $Z_{\lambda}^{u}(\phi)$ can be formally included in the general framework (1.1), where the function $f$ is seen as a tempered distribution, see Section 2.3, and $X=(Y,\nabla Y)$. It is then not surprising that the study of the nodal volume uses similar tools as those employed in the more general context of Section 1.1.
The CLT for the number of zeros on a growing interval of a stationary Gaussian process with $L^{2}$ covariance function and its second derivative was proved in [Cuz76] using the method of approximation by $m$-dependent processes, under the assumption of positivity of the limiting variance. This result was extended, under a similar square integrability condition, to the nodal length of a Gaussian field from $\mathbb{R}^{2}$ to $\mathbb{R}$ in [Kra01] using the method of Wiener chaos decomposition. The proof works in any dimension and was reproved under an additional smoothness assumption using the method of cumulants, see [Gas21t, Anc24]. In [Kra01], for the dimension $d=2$, a first general lower bound for the limiting variance appears, given by the variance of the second chaotic component, which for some constant $C$ is expressed as
$\operatorname{Var}(Z_{\lambda}^{(2))}(\phi)=C\int_{\mathbb{R}^{2}}\|\nabla^{2}r(x)\|^{2}-2d\langle\nabla r,\nabla r\rangle+d^{2}r(x)^{2}\mathrm{d}x=C\int_{\mathbb{R}^{2}}(\Delta r(x)+dr(x))^{2}\mathrm{d}x,$
where the equality follows from an application of the divergence theorem. The heuristic is valid in any dimension and proves that the limit is positive as long as $r$ is not an eigenfunction of the Laplace operator, which is never the case if $r\in L^{2}$.
For vector-valued random fields, i.e., when $k>1$, few results exist in the literature. The papers [Est16, Nic17, Aza24] prove a CLT for critical points and the Euler characteristic of random fields, and a positive lower bound is given under the assumption that the random field is isotropic or under the existence of a spectral density of the field near the origin. On the additional assumption of a smooth covariance function, normal asymptotics have been proved using the method of cumulants in the recent paper [Anc24]. We first prove a general theorem concerning the asymptotic normality of $Z_{\lambda}(\phi)$, based on a variance bound inherited from the strategy in [Gas21b] and the standard fourth moment theorem [Nua05] with the contraction principle. We define
$Z_{\lambda}^{u,(q+)}(\phi)=\sum_{k=q}^{+\infty}Z_{\lambda}^{u,(k)}(\phi),$
the $q$-tail of the variable $Z_{\lambda}^{u}(\phi)$. {mdframed}[backgroundcolor=white]
Theorem 1.5.
Let $q\geq 1$, and $\Omega\in L^{q}(V)$. There is a constant $C_{q}^{u}$ such that
$\lim_{\lambda\rightarrow+\infty}\operatorname{Var}(Z_{\lambda}^{u,(q+)}(\phi))\rightarrow C_{q}^{u},$
and
$Z_{\lambda}^{u,(q+)}(\phi)-\mathbb{E}[Z_{\lambda}^{(q+)}(\phi)]\underset{\lambda\rightarrow+\infty}{\longrightarrow}\mathcal{N}(0,C_{q}^{u}).$
Let us study first the projection on the first chaos of $Z_{\lambda}^{u}(\phi)$. By symmetry (see also Lemma 4.5), the first chaotic projection in the case $u=0$ cancels, so the next theorem is relevant for $u\neq 0$. {mdframed}[backgroundcolor=white]
Theorem 1.6.
Assume that spectral measure $\psi$ of $Y$ has a continuous density $\Sigma$ w.r.t the Lebesgue measure near the origin. Then
$\lim_{\lambda\rightarrow+\infty}\operatorname{Var}(Z_{\lambda}^{u,(1)}(\phi))=\Sigma(0)(u,u).$
If additionally, $r$ and its second differential are square integrable, and $\Sigma(0)(u,u)>0$ then $Z_{\lambda}^{u}(\phi)$ satisfies the CLT.
If the spectral measure $\psi$ does not admits a continuous density near the origin, one must use the general Corollary 1.2 to study the exact variance asymptotics of the first chaos. This theorem applies in particular if $\Omega$ is integrable. If the quantity $\Sigma(0)(u,u)$ cancels, one can use ideas similar to the next Theorem 1.7 (stated only for $u=0$) to study the variance of the second chaotic component, see also Remark 5.1.
In the case $u=0$, the two main obstructions in the case $k>1$ to prove general lower bound for the variance by an analysis of the second chaotic component, is that one cannot assume that the random vector $X(0)=(Y(0),\nabla Y(0))$ follows a standard Gaussian distribution by choosing adequate scalar products on $U$ and $V$ (see Remark 4.2), and that it is in general impossible, to compute explicitly the chaotic expansion of the nodal volume without any additional assumption on the field. The computation can be done in particular cases, see Lemma 4.5, under some assumptions of independence. The following theorem proves, under a square integrability property of the covariance function of $Y$, that the variance of the nodal volume converges to a positive constant under classical constraints on the field. {mdframed}[backgroundcolor=white]
Theorem 1.7.
Assume that $r$ and its second differential are square integrable on $V$, and one of the three following hypotheses.
  • (H1)
    There are bases of $U$ and $V$ such that $\operatorname{Var}(Y(0),\nabla Y(0))=\mathrm{Id}$
  • (H2)
    There is $u\in U$ such that for all $v\in u^{\perp}$, the process $\langle Y,u\rangle$ and $\langle Y,v\rangle$ are independent
  • (H3)
    $Y$ is a gradient field
Then there is a finite positive constant $C$ such that
$\lim_{\lambda\rightarrow+\infty}\operatorname{Var}(Z_{\lambda}(\phi))=C,$
and $Z_{\lambda}$ satisfies the CLT.
This theorem recovers and generalize many previously known results in the literature concerning the positivity of the lower bound for the nodal volume under the square integrability hypothesis.
– Hypothesis $(H1)$ :
This hypothesis assume a strong joint independence assumption on the random variable $(Y(0),\nabla Y(0))$, but assume nothing of the independence of coordinates processes. Following Remark 4.2, this hypothesis is always satisfied if $k=1$.
– Hypothesis $(H2)$ :
This second hypothesis treats the case of the volume of the intersection of the nodal set of a real random field with the nodal set of another independent vector field.
– Hypothesis $(H3)$ :
This last hypothesis proves the variance asymptotics and the CLT for the number of critical points of a random field without any additional assumptions. This hypothesis is independent of the first two because of the non-independence of the second derivatives. It recovers and generalizes some of the recent results concerning critical points such as [Nic17, Aza24].
We were not able to find a satisfying statement that encompasses the three situation, but the heuristic of proof for the positivity of the limiting variance, given in Section 5.2, is that in the three cases we are able to write the density $\Sigma$ of the spectral measure $\mu$ of the process $X=(Y,\nabla Y)$ as a sum of tensor products of two terms: one depending only on the spectral density of the underlying field $\sigma$, and the other one being a polynomial, that comes up from taking the Fourier transform of the derivatives. The conclusion follows by the simple geometric fact that a nonzero square integrable function cannot be supported on the zero set of a non-zero polynomial. In the case where $r\notin L^{2}(V)$, the situation is more intricate and is the content of the next Section.
1.3 Random waves and second chaos cancellation
We use the notation of the previous section. We define $\sigma$ to be the uniform probability measure on the sphere of radius $\sqrt{d}/2\pi$. Let $Y:V\rightarrow U$ be stationary Gaussian random process such that $\operatorname{Var}(Y(0),\nabla Y(0))$ is a standard Gaussian random vector. We say that $Y$ is a random wave if its spectral measure $\psi$ is supported on a sphere (necessary of radius $\sqrt{d}/2\pi$). Here we allow the dimension of $U$ to be greater than one, meaning that $Y$ can be seen as a collection of real random waves. Equivalently, $Y$ is a solution of the Laplace equation
$\Delta Y+dY=0.$
We say that $Y$ is a regular random wave if the spectral measure admits a continuously differentiable density w.r.t the measure $\sigma$. The particular case where $\psi=\mathrm{Id}\,\sigma$ (meaning that on an orthonormal basis, $Y$ is a collection of iid isotropic distributions), will be called the isotropic random wave model.
The isotropic random wave model has a long history. It was informally conjectured by Berry in [Ber77] to be the local limit of high-energy Laplace eigenfunctions on a generic manifold; see the introduction of [Gar23] for the heuristic of the conjecture, relevant literature on the topic, and various possible rigorous formulations of this conjecture. The conjecture implies, in particular, that the nodal set of an eigenfunction tends to equidistribute on the manifold, which supports another long-standing conjecture by Yau [Yau82] concerning the asymptotics of the nodal volume of high-energy eigenfunctions. To obtain a more tractable model, one can approximate an eigenfunction associated with $\lambda$ by a random sum of eigenfunctions with energy located in a small energy window $[\lambda-\varepsilon,\lambda+\varepsilon]$. This model is known as the monochromatic random wave model, and it has been shown that it serves as a good approximation to the isotropic Euclidean random wave model; see [Can16, Gas21, Die20].
On the torus and the sphere, particularly in dimensions $2$ and $3$, where the spectral decomposition of the Laplace operator is explicit, the model of monochromatic random waves and their nodal set has been intensively studied; see [Ros19, Mar23]. In particular, it has been observed that the variance of various local observables of the isotropic random wave (such as the nodal volume [Wig09, Rud08, Kri13, Mar16, Cam19, Mar20], the defect volume [Mar11], and the Euler characteristic [Cam18b]), in the high-energy limit, is of lower order than one might expect for a "generic" local observable, as suggested by Equation (4.1). This phenomenon is known as the Berry cancellation phenomenon and is purely a mathematical consequence of the second chaos cancellation as $\lambda$ approaches +$\infty$+ ∞. The proofs presented in the aforementioned papers exploit special properties of Bessel functions and the explicit chaos decomposition of the nodal volume functional to establish such general estimates. In the following, we restrict our study to the Euclidean case, which represents the scaling limit of monochromatic random waves on a generic compact manifold. It is expected that this cancellation phenomenon also appears in a broader Riemannian context; see the forthcoming paper [Ste25].
We first prove that this second chaos cancellation can be understood in a simple manner from a spectral point of view by the more general Theorem 1.4, shedding new light on this cancellation phenomenon, which is valid for a more general class of regular random waves. For instance, the second identity in Equation (1.1.2), which heuristically explains the cancellation phenomenon from the spatial representation of random waves, becomes trivial in the Fourier domain and reduces to the identity
$\|\xi\|^{4}-2d\|\xi\|^{2}+d^{2}=(-\|\xi\|^{2}+d)^{2}=0\quad\text{if}\quad\|\xi\|=\sqrt{d},$
which was actually the source of motivation to start the writing of this paper in the first place. {mdframed}[backgroundcolor=white]
Theorem 1.8.
Assume that there are basis of $U$ and $V$ such that $\operatorname{Var}(Y(0),\nabla Y(0))=\mathrm{Id}$. If $Y$ is not a random wave then
$\liminf_{\lambda\rightarrow+\infty}\operatorname{Var}(Z_{\lambda}^{(2)}(\phi))>0.$
Conversely, if $Y$ is a regular random wave then
$\lim_{\lambda\rightarrow+\infty}\operatorname{Var}(Z_{\lambda}^{(2)}(\phi))=0.$
In this toy-model example, one explicitly computes the variance of the second chaotic projection to provide a partial equivalence between the inclusion of the support of the spectral measure in the sphere and the second chaos cancellation. In more complicated settings, such as those presented in Theorem 1.8, where explicit computation is not always feasible, one can still establish the positivity of the liminf if the support of the spectral measure is not contained within the zero set of a quadratic form. This approach forms the basis of the proof of Theorem 1.7, under the additional square-integrability assumption. We conjecture that some version of the equivalence given by Theorem 1.8 holds true for critical points, but the computation of the second chaotic component in this context proves to be challenging.
The heuristic, valid for "isotropic" observables such as the nodal volume or some of the aforementioned observables, implies that the isotropic cone of the second chaotic component must form a sphere, whose radius needs to be determined. In the case of the nodal volume, this sphere has a radius precisely $\sqrt{d}/2\pi$. The image of the spectral measure of the field $X=(Y,\nabla Y)$ is a sphere of radius $\sqrt{d}/2\pi$ if and only if $Y$ is a random wave, and the conclusion follows from the second point of Theorem 1.4.
This theorem is a multidimensional generalization of some of the findings in [Lac21], where it was proved in the case $V=\mathbb{R}$ that the limiting variance is positive (possibly infinite) if and only if the process is not a cosine with a Gaussian amplitude. In fact, if the spectral measure, restricted to a domain at a positive distance from the sphere of radius $\sqrt{d}/2\pi$, is not square-integrable, then a straightforward byproduct of the proof, along with Lemma 3.8 and Lemma 3.9, shows that the variance diverges to +$\infty$+ ∞. Together with the discussion after Theorem 1.4, this situation can also lead to a CLT for the quantity $Z_{\lambda}$, inherited from a CLT on $Z_{\lambda}^{(2)}$, provided one can quantify the singularity of the measure $\psi*\psi$ near the origin.
We recover and generalize, for any dimension $d$ and $k$, several of the aforementioned results concerning the second chaos cancellation of the volume of the nodal intersection of $k$ isotropic and independent random waves in $\mathbb{R}^{d}$. For more general $u$-levels, we refer to Remark 5.1. To understand the variance asymptotics of $Z_{\lambda}(\phi)$, one must examine the $4$-th chaos, which is the content of the next Section 1.4.
1.4 Asymptotics of the fourth chaos
We showed in the previous paragraph that regular random waves, and in particular isotropic random waves, are subject to the general cancellation phenomenon of Theorem 1.4. In this section, we derive the exact variance asymptotics of the nodal volume of the isotropic random wave model, following the heuristic outlined after Theorem 1.8. The result we obtain encompasses and generalizes many findings in the literature concerning such asymptotics, which have primarily focused on dimensions $2$ and $3$.
The case $d=2$ is well-studied, with significant results found in works such as [Cam19, Kri13, Nou19] across various settings (spherical harmonics, arithmetic random waves, planar random waves). The exact asymptotics of the variance of the nodal length have been explicitly computed : there exists an explicit positive constant $C$ such that
$\lim_{\lambda\rightarrow+\infty}\frac{\operatorname{Var}(Z_{\lambda}(\mathds{1}_{B(0,1)}))}{\log(\lambda)}=\lim_{\lambda\rightarrow+\infty}\frac{\operatorname{Var}(Z_{\lambda}^{(4)}(\mathds{1}_{B(0,1)}))}{\log(\lambda)}=C,$
and $Z_{\lambda}$ satisfies a CLT. When $d=3$, it has been proved in [Dal19] the existence of a positive constant $C$ such that
$\lim_{\lambda\rightarrow+\infty}\operatorname{Var}(Z_{\lambda}(\mathds{1}_{B(0,1)}))=C,$
as well as a CLT. The next theorem proves a general version of these two facts. {mdframed}[backgroundcolor=white]
Theorem 1.9.
Let $Y$ be the isotropic random wave model defined in introduction. Then there is a positive constant $C_{d}$ such that as $\lambda$ goes to +$\infty$+ ∞
$\operatorname{Var}(Z_{\lambda}(\phi))\simeq\left\{\begin{array}[]{llll}C_{2}\log(\lambda)\;\text{ if }d=2,\\ C_{d}\;\qquad\;\;\,\text{ if }d\geq 3,\\ \end{array}\right.$
and $Z_{\lambda}(\phi)$ satisfies the CLT.
The heuristic that leads to the distinction between $d=2$ and $d>2$ is as follows, and can already be seen from the general discussion that follows Theorem 1.4, and corroborates the heuristics for the fluctuations of the polyspectra elaborated in [Gro24]. From the discussion in the previous Section 1.3, we know that the variance of the second chaotic component cancels, despite having a covariance function that is not square-integrable. To obtain the exact asymptotics of the variance, one must examine the next non-zero chaotic component, namely the fourth one. To compute the fourth chaos and understand the difference between the cases $d=2$ and $d\geq 3$, we need to analyze, in light of Theorem 1.1, the behavior of the four-fold convolution of the uniform measure on the sphere $\sigma$ near 00.
– If $d=2$:
In this case, one can show that $\sigma*\sigma$ is not square-integrable, or equivalently, that $\Omega\notin L^{4}(V)$. This means that $\sigma^{*4}$ diverges near the origin. Its exact asymptotics are computed in Lemma 4.8, where it is shown to behave as $|\log(x)|$ near the origin. The exact asymptotics of the variance of the fourth chaos then follow from the principles underlying Theorem 1.4, proving that the limit is non-degenerate. As for the higher chaotic components, note that $\Omega\in L^{6}(V)$. Their variances converge to a constant, and they do not contribute to the asymptotics of the total variance, which is of order $\log(\lambda)$. In other words, $Z_{\lambda}$ is asymptotically equal to its fourth chaotic component $Z_{\lambda}^{(4)}$.
– If $d\geq 3$:
In this case, $\sigma*\sigma$ is square-integrable, or equivalently, $\Omega\in L^{4}(V)$. This means that $\sigma^{*4}(0)=\|\sigma*\sigma\|_{2}^{2}$ is well-defined and positive: the variance of the fourth chaos converges to a constant, and we will prove, by the same method used in the case $d=2$, that this constant is positive. The higher chaotic components will also contribute a positive term to the total variance, and unlike the case $d=2$, there is no dominance of a single chaotic component.
{mdframed}
[backgroundcolor=white]
Corollary 1.10.
Assume that $d\geq 2$ and that the process $Y:V\rightarrow\mathbb{R}$ has an isotropic distribution. Then
$\liminf_{\lambda\rightarrow+\infty}\operatorname{Var}(Z_{\lambda}(\phi))>0.$
This corollary is a straightforward consequence of Theorem 1.8 and Theorem 1.9. It implies that there cannot be any variance cancellation phenomenon under the assumption of isotropy as soon as $d\geq 2$. In dimension $d=1$ the cosine process with random amplitude is the only counterexample, as asserts [Lac21].
2 Hermite polynomials and Wiener chaos expansion
2.1 Hermite polynomials
In this section, we expose some material about Wiener chaos expansion on a finite dimensional Gaussian space, see [Pec11, Nou12] for general references. The following material is standard but exposed in a way that avoids the use of coordinates to obtain compact and elegant formulas for chaos projections and covariances, in term of the natural scalar product on the space of multilinear maps induced by the underlying Gaussian measure.
Let $W$ be a finite dimensional real vector space of dimension $n$ and $W^{*}$ be its dual space. For a integer $q\geq 0$ we denote $(W^{*})^{\otimes q}$ the space of $q$-linear maps on $W$. Its subspace of symmetric $q$-linear maps can naturally be identified with the space of homogeneous polynomial of degree $q$ on $W$ via the isomorphism $f\mapsto(w\mapsto f(w,\ldots,w))$. In what follows, we will simply write $f(w)=f(w,\ldots,w)$ when $f$ is a symmetric $q$-linear map.
[…]
$\langle\ell_{1}\otimes\ldots\otimes\ell_{q},m_{1}\otimes\ldots\otimes m_{q}\rangle=\prod_{k=1}^{q}\langle\ell_{k},m_{k}\rangle.$
It generalizes the Frobenius scalar product for matrices (i.e. the case $q=2$). We will denote $w^{*}$ the linear form $\langle w,\,.\,\rangle$. For a vector $w\in W$, the Radon–Nikodym derivative of the translated measure $\eta(.\,-w)$ with respect to $\eta$ is given by
$\frac{\mathrm{d}\eta(.\,-w)}{\mathrm{d}\eta}:x\longmapsto\exp\left(\langle w,x\rangle-\frac{1}{2}\|w\|^{2}\right).$
It allows us to define for $x\in W$ the symmetric $q$-linear Hermite form $H_{x}^{q}$ via the Taylor expansion
$\frac{\mathrm{d}\eta(.\,-w)}{\mathrm{d}\eta}:x\mapsto\sum_{q=0}^{+\infty}\frac{1}{q!}H_{x}^{q}(w),\quad\text{with}\quad H_{x}^{q}=(-1)^{q}\exp\left(\frac{1}{2}\|x\|^{2}\right)D^{q}_{x}\left[\exp\left(-\frac{1}{2}\|\,.\,\|^{2}\right)\right].$
[…]
$E(\eta,\tau)=\left\{\Omega\in(W^{*}\otimes U^{*})^{*}\leavevmode\nobreak\ \middle|\leavevmode\nobreak\ \forall\ell\in W^{*},\;\;m\in U^{*},\quad\|\ell\|^{2}+\|m\|^{2}+2\Omega(\ell,m)\geq 0\right\},$
and $E^{*}(\eta,\tau)$ the interior of $E(\eta,\tau)$, i.e. where the inequality in the definition is strict. When $W=U$ and $\eta=\tau$ we simply denote them $E(\eta)$ and $E^{*}(\eta)$. For $\Omega\in E(\eta,\tau)$, we define the Gaussian measure $\eta\otimes_{\Omega}\tau$ on $W\times U$ so that its marginals are given by $\eta$ and $\tau$ respectively, and such that $\Omega$ is the covariance bilinear map between the two marginals. For instance, the choice of $\Omega=0$ leads to the product measure on $W\times U$. One has following Gaussian integral, for $w\in W$ and $u\in U$
$\int_{W\times U}\exp\left(\langle w,x\rangle-\frac{1}{2}\|w\|^{2}\right)\left(\langle u,y\rangle-\frac{1}{2}\|u\|^{2}\right)\mathrm{d}\eta\!\otimes_{\Omega}\!\tau(x,y)=\exp\left(\Omega(w^{*},u^{*})\right).$
Developing inside the integral on the left using Equation (2.1), and using the Taylor expansion of the exponential of the right, one get by identifying the $q$-linear maps for different $q$, the relations for $q,q^{\prime}\geq 0$
$\int_{W\times U}H_{x}^{q}(w)H_{y}^{q^{\prime}}(u)\mathrm{d}\eta\otimes_{\Omega}\tau(x,y)=\delta_{qq^{\prime}}q!\Omega(w,u)^{q}=\delta_{qq^{\prime}}q!\Omega^{\otimes{q}}((w^{*})^{\otimes q},(u^{*})^{\otimes q})),$
where $\Omega^{\otimes q}$ is seen as a bilinear map on $(W^{*})^{\otimes q}\times(U^{*})^{\otimes q}$.
2.2 Wiener chaos expansion
From now on, we assume that $W=U$ and $\eta=\tau$, so that $\Omega$ is a bilinear map on $W^{*}$. The choice of $\Omega=\langle\,.\,,\,.\,\rangle$ in the previous paragraph leads to $\eta\otimes_{\Omega}\eta$ being the diagonal measure on $W\times W$ induced by $\eta$. In that case, Equation (2.1) yields the classical orthogonality relations
$\int_{W}H_{x}^{q}(w)H_{x}^{q^{\prime}}(u)\mathrm{d}\eta(x)=\delta_{qq^{\prime}}q!\langle w,u\rangle^{q}.$
This relation allows us to detail the classical Wiener chaos decomposition for the space $L^{2}(\eta)$ in our framework. We define the $q$-Hermite chaos spaces of functions
$\mathcal{H}_{q}=\mathrm{Vect}\left(\left(x\mapsto H^{q}_{x}(w)\right)_{w\in W}\right).$
The previous orthogonality relation (2.2) implies that the vector spaces $\mathcal{H}_{q}$ and $\mathcal{H}_{q^{\prime}}$ are orthogonal for different $q,q^{\prime}$. Moreover, their union coincides with the set of polynomials on $W$ and is thus dense in $L^{2}(\eta)$. We directly deduce the orthogonal decomposition
$L^{2}(\eta)=\overline{\bigoplus_{q=0}^{\infty}\mathcal{H}_{q}}.$
For $f\in L^{2}(\eta)$, we define the symmetric $q$-linear map $f_{q}$ on $W$ and the $q$th Hermite projection $\pi_{q}(f)$ as
$f_{q}=\int_{W}f(x)H_{x}^{q}\mathrm{d}\eta(x)\quad\text{and}\quad\pi_{q}(f):x\mapsto\frac{1}{q!}\langle f_{q},H^{q}_{x}\rangle.$
Then for $q,q^{\prime}\in\mathbb{N}$ and $w\in W$, one has
$\pi_{q}(H_{.}^{q^{\prime}}(w))(y)=\frac{1}{q!}\left\langle\int_{W}H_{x}^{q^{\prime}}(w)H_{x}^{q}\mathrm{d}\eta(x)\,,\,H_{y}^{q}\right\rangle=\delta_{qq^{\prime}}\left\langle(w^{*})^{\otimes q},H_{y}^{q}\right\rangle=\delta_{qq^{\prime}}H_{y}^{q}(w),$
where the second equality follows from Equation (2.2). It shows that the operator $\pi_{q}$ is indeed the orthogonal projection of $L^{2}(\eta)$ onto the $q$-chaos space $\mathcal{H}_{q}$. If we go back to a general coupling $\Omega$ between $\eta$ and itself on $W^{*}\times W^{*}$, one obtain, for $f,g\in L^{2}(\eta)$, the following decomposition
$\int_{W\times W}f(x)g(y)\mathrm{d}\eta\otimes_{\Omega}\eta(x,y)$ $=\sum_{q,q^{\prime}=0}^{+\infty}\int_{W\times W}\pi_{q}f(x)\pi_{q^{\prime}}g(y)\mathrm{d}\eta\otimes_{\Omega}\eta(x,y)$
$=\sum_{qq^{\prime}=0}^{+\infty}\frac{1}{q!q^{\prime}!}\int_{W\times W}\langle f_{q},H_{x}^{q}\rangle\langle g^{q^{\prime}},H_{x}^{q^{\prime}}\rangle\mathrm{d}\eta\otimes_{\Omega}\eta(x,y)$
$=\sum_{q=0}^{+\infty}\frac{1}{q!}\Omega^{\otimes q}(f_{q},g_{q}).$ (7)
2.3 Hermite expansion of distributions
Let $f$ be a tempered distribution on $W$, and $(f_{q})_{q}$ as sequence of $q$-linear forms such that it holds in the sense of tempered distribution the equality
$f=\sum_{q=0}^{+\infty}\langle f_{q},H_{x}^{q}\rangle.$
Note by a density argument, one necessary has
$f_{q}=f(h_{q}),\quad\text{with}\quad h_{q}:x\mapsto H_{x}^{q}\exp\left(-\frac{\|x\|^{2}}{2}\right).$
For $\Omega\in E^{*}(\eta)$, the density $\rho_{\Omega}$ of the measure $\eta\otimes_{\Omega}\eta$ w.r.t the Lebesgue measure on $W\times W$ is well-defined and is a function in the Schwartz class. Then the mapping $C_{f}:E^{*}(\eta)\rightarrow\mathbb{R}$ given by
$C_{f}:\Omega\mapsto(f\otimes f)(\rho_{\Omega})$
is well defined, and as before it holds that for $\Omega\in E^{*}(\eta)$,
$C_{f}(\Omega)=\sum_{q=0}^{+\infty}\frac{1}{q!}\Omega^{\otimes q}(f_{q},f_{q}).$
It shows that the function $C_{f}$ is analytic on $E^{*}(\eta)$, and the right-hand side corresponds to its series expansion. Note that if $f\in L^{2}(\eta)$, the function $C_{f}$ can be continuously extended to $E(\eta)$. Let $A$ be a compact set of $E^{*}(\eta)$ and $\Omega\in A$. The integral remainder in Taylor expansion implies that
$\left|C_{f}(\Omega)-\sum_{k=0}^{q-1}\Omega^{\otimes k}(f_{k},f_{k})\right|\leq\left(\sup_{\Omega\in A}\frac{1}{q!}\|D^{q}C_{f}\|\right)\|\Omega\|^{q}.$
This expression is an alternative to Arcones inequality, see [Arc94], which has the advantage of being valid for any tempered distribution $f$. This bound is inspired from the results in [Gas21b], and holds without any square integrability assumption on $f$, the downside being that one must restrict $\Omega$ to $E^{*}(\eta)$, otherwise $C_{f}$ might be ill-defined (a typical example is the choice of function $f=\delta_{0}$).
3 Positive definite functions, Hermitian form and convolution
The Fourier transform of a covariance function of a process $Y:V\rightarrow W$ is a measure taking values in the space of Hermitian forms on $W^{*}$. In the following, we expose the basic theory of such measure.
3.1 Positive Hermitian forms
Let $W$ be a finite dimensional complex vector space. We denote by $\overline{W}$ the conjugate vector space of $W$ and $\operatorname{Sesq}(W)=W^{*}\otimes\overline{W}^{*}$, the space of sesquilinear forms on $W^{*}$, which is naturally identified with the space of linear maps between $\overline{W}$ and $W^{*}$. The image of an element $\Sigma\in\operatorname{Sesq}(W)$, denoted by $\operatorname{Im}(\Sigma)$, is the image of this mapping, as a subspace of $W^{*}$. If $\Sigma\in\operatorname{Sesq}(W)$, we define $\Sigma^{c}\in\operatorname{Sesq}(W)$, the complex conjugate of $\Sigma$, via the formula for $w_{1},w_{2}\in W$
$\Sigma^{*}(w_{1},w_{2})=\overline{\Sigma(w_{2},w_{1})}.$
Let $\mathcal{H}(W)$ be the space of Hermitian forms on $W$, that is the subspace of sesquilinear forms on $W$ invariant by complex conjugate. We simply write $\Sigma(w)=\Sigma(w,w)$ when $\Sigma$ is a Hermitian form, since it depends only on the diagonal evaluations via the polarization identity.
$\Sigma(w_{1},w_{2})=\frac{1}{4}\left(\Sigma(w_{1}+w_{2})-\Sigma(w_{1}-w_{2})+i\Sigma(w_{1}+iw_{2})-i\Sigma(w_{1}-iw_{2})\right).$
An Hermitian form $\Sigma$ is positive semi-definite if additionally, for all $w\in W$, $\Sigma(w)\geq 0$. We will denote by $\mathcal{H}^{+}(W)$ the set of positive semi-definite Hermitian form on $W$.
We now prove a few technical lemmas concerning Hermitian forms. Let $\Sigma_{1},\ldots,\Sigma_{q}$ Hermitian forms on complex vector spaces $W_{1},\ldots,W_{q}$. We define the Hermitian form on $W_{1}\otimes\ldots\otimes W_{q}$
$\bigotimes_{k=1}^{q}\Sigma_{k}:(w_{1}\otimes\ldots\otimes w_{q})\mapsto\prod_{k=1}^{q}\Sigma_{k}(w_{k}).$
In particular, if $\Sigma$ is a Hermitian form on $W$, then $\Sigma^{\otimes q}$ is a Hermitian form on $W^{\otimes q}$. {mdframed}[backgroundcolor=white]
Lemma 3.1.
Let $\Sigma_{1},\ldots,\Sigma_{q}$ positive semi-definite Hermitian forms on complex vector spaces $W_{1},\ldots,W_{q}$. Then $\bigotimes_{k=1}^{q}\Sigma_{k}$ is a positive semi-definite Hermitian form on $W_{1}\otimes\ldots\otimes W_{q}$.
Proof.
By the diagonalization theorem for Hermitian forms, for $1\leq k\leq q$ there is a basis of $W_{k}$, denoted $w^{1}_{k},\ldots,w^{d_{k}}_{k}$ such that $\Sigma_{k}(w^{i}_{k},w^{j}_{k})=0$ if $i\neq j$. Let $w\in\bigotimes_{k=1}^{q}W_{k}$. Decomposing $w$ on the induced basis on $\bigotimes_{k=1}^{q}W_{k}$, so that
$w=\sum_{i_{1},\ldots,i_{q}=1}^{d_{i_{1}},\ldots,d_{i_{q}}}\alpha_{k}^{i_{k}}\bigotimes_{k=1}^{q}w_{k}^{i_{k}}$
one gets
$\bigotimes_{k=1}^{q}\Sigma_{k}(w)=\sum_{i_{1},\ldots,i_{q}=1}^{d_{i_{1}},\ldots,d_{i_{q}}}\prod_{k=1}^{q}|\alpha_{k}^{i_{k}}|^{2}\Sigma_{k}(w_{k}^{i_{k}})\geq 0,$
which proves that the measure $\mu_{1}\otimes\ldots\otimes\mu_{q}$ is positive. ∎
{mdframed}
[backgroundcolor=white]
Lemma 3.2.
[…]
$\Sigma^{\otimes 2}(f,f)=\|f(P(\cdot),P(\cdot))\|^{2}.$
Proof.
Fixing a basis on $W$ this identity simply reads
$\operatorname{Tr}(\bar{P}P^{T}fP\bar{P}^{T}\bar{f})=\|P^{T}fP\|^{2}.$
Given a symmetric element $f\in W^{*}\otimes W^{*}$, we define its isotropic cone $C(f)$ as the set
$C(f)=\left\{w\in W\leavevmode\nobreak\ \middle|\leavevmode\nobreak\ f(w)=0\right\}.$
{mdframed}
[backgroundcolor=white]
Lemma 3.3.
Let $\Sigma\in\mathcal{H}^{+}(W^{*})$ and $f\in W\otimes W$. Then
$\Sigma^{\otimes 2}(f,f)=0\Longleftrightarrow\operatorname{Im}(\Sigma)\subset C(f).$
Proof.
[…]
$\Sigma(\cdot,\cdot)=\langle\sqrt{\Sigma}(\cdot),\sqrt{\Sigma}(\cdot)\rangle,$
From the previous lemma,
$\Sigma^{\otimes 2}(f,f)=\left\|f\left(\sqrt{\Sigma}(\,.\,),\sqrt{\Sigma}(\,.\,)\right)\right\|^{2}.$
Since the maps $\Sigma$ and $\sqrt{\Sigma}$ have the same image, the right hand side cancels if and only if $\operatorname{Im}(\Sigma)\subset C(f)$. ∎
3.2 Positive Hermitian measure and Bochner Theorem
Let $V$ be a finite dimensional real vector space and $\mathcal{B}(V)$ be its Borel $\sigma$-field. Let $W$ be a finite dimensional complex vector space. A map $\mu:\mathcal{B}(V)\rightarrow\operatorname{Sesq}(W)$ is a Hermitian measure over $W$ if for all $w\in W$, the mapping $\mu_{w}:E\mapsto\mu(E)(w)$ is a bounded signed measure on $(V,\mathcal{B}(V))$, and is positive if it is a non-negative measure. If one fix a basis of $W$, then on can think of $\mu$ as a matrix of complex-valued measures such that for all measurable set $E$, such that the quantity $\mu(E)$ is a Hermitian matrix, positive semi-definite if $\mu$ is positive.
A function $\Omega:V\rightarrow\operatorname{Sesq}(W)$ is positive semi-definite if for all choice of vectors $v_{1},\ldots,v_{m}\in V$ and $w_{1},\ldots,w_{m}\in W$, one has
$\sum_{i,j=1}^{m}\Omega(v_{i}-v_{j})(w_{i},w_{j})\in\mathbb{R}_{+}.$
When $W=\mathbb{C}$, this is the usual definition of positive semi-definite functions. Bochner theorem then states that there is an equivalence between being a continuous positive semi-definite complex valued function, and being the Fourier transform of a bounded measure on the real line. We show that this theorem has a straightforward extension to functions taking values in $\mathcal{S}q(W)$. {mdframed}[backgroundcolor=white]
Lemma 3.4.
Let $\Omega:V\rightarrow\mathcal{S}q(W)$. The two propositions are equivalents.
  • (1)
    $\Omega$ is a continuous positive semi-definite function.
  • (2)
    $\Omega$ is the Fourier transform of a positive Hermitian measure $\mu$ on $V^{*}$ over $W$.
Proof.
Let us prove first the direct implication. For $w\in W$, the function $\Omega_{w}:v\mapsto\Omega(v)(w,w)$ is a continuous complex valued positive semi-definite function. By the usual Bochner Theorem, $\Omega_{w}$ is the Fourier transform of a non-negative bounded measure on $V^{*}$, denoted $\mu_{w}$. Let $E$ be a measurable set in $V^{*}$. The linearity of the Fourier transform and the polarization identity (3.1) imply that the mapping
$\mu(E):w\mapsto\mu_{w}(E)$
defines a positive semi-definite Hermitian form on $W$ and that $\hat{\mu}=\Omega$. Conversely, Assume that $\Omega=\hat{\mu}$ is a positive Hermitian measure on $V^{*}$ over $W$. Since the measure $\mu$ is bounded, the function $\Omega$ is continuous. Let $v_{1},\ldots,v_{m}\in V$ and $w_{1},\ldots,w_{m}\in W$. Then
$0\leq\int_{V^{*}}\mathrm{d}\mu(\xi)\left(\sum_{i=1}^{m}w_{i}e^{2i\pi\xi(v_{i})}\right)=\sum_{i,j=1}^{m}\int_{V^{*}}e^{2i\pi\xi(x_{i}-x_{j})}\mathrm{d}\mu(w_{i},w_{j})=\sum_{i,j=1}^{m}\Omega(x_{i}-x_{j})(w_{i},w_{j}).$
This shows that $\Omega$ is a positive semi-definite function.∎
Remark 3.5.
If $W$ is a real vector space instead of a complex vector space, we can embed it in its complexification $W_{\mathbb{C}}=W\otimes\mathbb{C}$ via the mapping $w\mapsto w\otimes 1$. A map $\Omega:V\rightarrow(W^{*})^{\otimes 2}$ can be seen as a map taking values in $\operatorname{Sesq}(W_{\mathbb{C}})$ via this embedding, and one can extend the notion of positive definiteness for such bilinear maps $\Omega$. The conclusion of Lemma 3.4 then holds if one replace $(2)$ by saying that $\Omega$ is the Fourier transform of a (symmetric) positive Hermitian measure over $W_{\mathbb{C}}$. In particular, if $X:V\rightarrow W$ is a stationary Gaussian process with continuous covariance function $\Omega$, then its spectral measure $\mu=\widehat{\Omega}$ is a symmetric positive Hermitian measure over $W_{\mathbb{C}}$.
3.3 Convolution of Hermitian measures
Let $q\geq 1$, $W_{1},\ldots,W_{q}$ be complex vector spaces and $\mu_{1},\ldots,\mu_{q}$ be Hermitian measures over $W_{1},\ldots,W_{q}$ respectively. We define the Hermitian measure $\mu_{1}*\ldots*\mu_{q}$ on $V$ over $\bigotimes_{k=1}^{q}W_{k}$ by the formula, for all vectors $w_{1},\ldots,w_{q}$ in $W_{1},\ldots,W_{q}$ respectively, by the formula
$\mu_{1}*\ldots*\mu_{q}(w_{1}\otimes\ldots\otimes w_{q})=\mu_{1}(w_{1})*\ldots*\mu_{q}(w_{q}),$
where on the right, the symbol * denotes the usual convolution of measures. Alternatively, one can construct the measure $\mu_{1}*\ldots*\mu_{q}$ as the push-forward of the Hermitian measure $\mu_{1}\otimes\ldots\otimes\mu_{q}$ on $V^{q}$ by the addition map from $V^{q}$ to $V$, meaning that for a measurable subset $E$ of $V$ one has
$\mu_{1}*\ldots*\mu_{q}(E)=\int_{V^{q}}\mathds{1}_{E}(x_{1}+\ldots+x_{q})\mathrm{d}(\mu_{1}\otimes\ldots\otimes\mu_{q})(x_{1},\ldots,x_{q}).$
If $\mu$ is a measure on $V$ taking value in the space of Hermitian forms on a complex space $W$, we define its $q$-fold convolution $\mu^{*q}$ as the convolution with itself $q$ times. {mdframed}[backgroundcolor=white]
Lemma 3.6.
Assume that for $1\leq k\leq q$, the measures $\mu_{k}$ are positives Hermitian measures over $W_{1},\ldots,W_{q}$ respectively. Then the measures $\mu_{1}\otimes\ldots\otimes\mu_{q}$ and $\mu_{1}*\ldots*\mu_{q}$ are positive Hermitian measures.
Proof.
The positivity of $\mu_{1}\otimes\ldots\otimes\mu_{q}$ follows directly from Lemma 3.1. The positivity of $\mu_{1}*\ldots*\mu_{q}$ then follows from the definition of convolution. ∎
Let $\nu$ be a bounded measure on $V$. We say that a positive Hermitian measure $\mu$ is absolutely continuous w.r.t $\nu$ if for all $w\in W$, the measure $\mu_{w}$ is absolutely continuous w.r.t to $\nu$, and we write it $\mu\ll\nu$. For instance, if $W$ comes equipped with an Hermitian scalar product (so one can take the trace of a sesquilinear form on $W$), then one can define the trace measure $\mu^{\operatorname{Tr}}:E\mapsto\operatorname{Tr}(\mu(E))$, and it is straightforward to see that $\mu\ll\mu^{\operatorname{Tr}}$. By the Radon-Nikodym theorem, there exists a non-negative function $\Sigma:V\rightarrow\mathcal{S}q(W)$, uniquely defined up to a $\nu$-null set, such that $\mathrm{d}\mu=\Sigma\,\mathrm{d}\nu$. Such a decomposition of $\mu$ will be called a representation of $\mu$. Note that for such a representation, one always has $\mu^{\operatorname{Tr}}\ll\nu$.
Given a representation of $\mu$, we define the image of $\mu$, denoted by $\operatorname{Im}(\mu)$, as the smallest closed subset $A\subset W^{*}$ such that
$\nu\left(\left\{x\in V\leavevmode\nobreak\ \middle|\leavevmode\nobreak\ \operatorname{Im}(\Sigma(x))\not\subset A\right\}\right)=0.$
This notion coincides with the essential range of the measurable function $(x,w)\mapsto\Sigma(x)(w,.)$, and is independent of the representation of $\mu$. {mdframed}[backgroundcolor=white]
Lemma 3.7.
The image of a positive Hermitian measure $\mu$ does not depend on a choice of representation.
Proof.
Let $\rho$ be another measure such that $\mu\ll\rho$. Up to replacing $\rho$ by $\nu+\rho$, and symmetry, we can assume that $\nu\ll\rho$. Let $\mathrm{d}\mu=\Gamma\mathrm{d}\rho$ and $\mathrm{d}\nu=\alpha\,\mathrm{d}\rho$. By the chain rule, $\Gamma=\alpha\Sigma$, so that $\Gamma$ and $\Sigma$ have the same images on the subset $\{\alpha>0\}$. The conclusion follows, since $0\in A$ and the measures $\nu$ and $\rho$ share the same null sets on the subset $\{\alpha>0\}$. ∎
3.4 Hermitian measure and approximation of unity
Let $\gamma=|\hat{\phi}|^{2}$, where $\phi$ is a test function as defined in introduction, and recall, for a measure $\nu$, the definitions of $D^{-}\nu$, $D^{+}\nu$ and $D\nu$. The proof of the following lemma is inspired from [Sae96].
{mdframed}
[backgroundcolor=white]
Lemma 3.8.
Let $\nu$ be a bounded measure on $V$. Then
$\liminf_{\lambda\rightarrow+\infty}\int_{V}\gamma_{\lambda}(x)\mathrm{d}\nu(x)\geq\|\gamma^{-}\|_{1}D\nu^{-}(0),$
and
$\limsup_{\lambda\rightarrow+\infty}\int_{V}\gamma_{\lambda}(x)\mathrm{d}\nu(x)\leq\|\gamma^{+}\|_{1}D\nu^{+}(0).$
Proof.
One can assume that $\gamma$ is radially decreasing, since one has the bounds $\gamma^{-}\leq\gamma\leq\gamma^{+}$. One has
$\int_{V}\gamma_{\lambda}(x)\mathrm{d}x=\int_{0}^{+\infty}\operatorname{Vol}\{\gamma_{\lambda}>t\}\mathrm{d}t,$
meaning that the function $t\rightarrow\operatorname{Vol}\{\gamma_{\lambda}>t\}$ is integrable on $\mathbb{R}_{+}$. More generally
$\int_{V}\gamma_{\lambda}(x)\mathrm{d}\nu(x)$ $=\int_{V}\int_{0}^{\gamma_{\lambda}(x)}\mathrm{d}t\mathrm{d}\nu(x)$
$=\int_{0}^{+\infty}\nu\{\gamma_{\lambda}>t\}\mathrm{d}t$
$=\int_{0}^{+\infty}\frac{\nu(\frac{1}{\lambda}\{\gamma>t\})}{\lambda^{d}}\mathrm{d}t$
$=\int_{0}^{+\infty}\frac{\nu(\frac{1}{\lambda}\{\gamma>t\})}{\operatorname{Vol}(\frac{1}{\lambda}\{\gamma>t\})}\operatorname{Vol}\{\gamma>t\}\mathrm{d}t$
The set $\{\gamma>t\}$ is a centered ball, whose radius is decreasing and finite as soon as $t>0$. For $\varepsilon>0$, there is a constant $C_{\varepsilon}$ such that
$\int_{V}\gamma_{\lambda}(x)\mathrm{d}\nu(x)\geq\inf_{R\leq\frac{C_{\varepsilon}}{\lambda}}\frac{\nu(B(0,R))}{\operatorname{Vol}(B(0,R))}\int_{\varepsilon}^{+\infty}\operatorname{Vol}\{\gamma>t\}\mathrm{d}t.$
[…]
For a non-negative bounded measure $\nu$ on $V$, we define $\|\nu\|_{2}=\|f\|_{2}$ if the measure $\nu$ has a density $f\in L^{2}(V)$ w.r.t the Lebesgue measure, and $\|\nu\|_{2}=+\infty$ otherwise. Alternatively, by Plancherel theorem, $\|\nu\|_{2}=\|\hat{\nu}\|_{2}$, where $\hat{\nu}$ is the Fourier transform of $\nu$, which is a continuous bounded function on $V^{*}$. {mdframed}[backgroundcolor=white]
Lemma 3.9.
Let $\nu$ be a non-zero symmetric bounded measure on $V$. Then
$D\nu*\nu(0)=\|\nu\|_{2}^{2}.$
Proof.
Let $h$ be the Fourier transform of the function $\frac{\mathds{1}_{B(0,1)}}{\operatorname{Vol}(B(0,1))}$. Then $|h(x)|\leq h(0)=1$, $h$ is continuous, and
$\frac{\nu*\nu(B(0,R))}{\operatorname{Vol}(B(0,R))}=\int_{V^{*}}h\left(\frac{x}{R}\right)\hat{\nu}^{2}(x)\mathrm{d}x.$
If $\|\nu\|_{2}<+\infty$ then this quantity converges towards $\|\nu^{2}\|_{2}^{2}$ by dominated convergence and Plancherel theorem. If $\|\nu\|_{2}=+\infty$, one has the lower bound
$\frac{\mathds{1}_{B(0,R)}}{\operatorname{Vol}(B(0,R))}\geq\frac{1}{2^{d}}\frac{\mathds{1}_{B(0,R/2)}}{\operatorname{Vol}(B(0,R/2))}*\frac{\mathds{1}_{B(0,R/2)}}{\operatorname{Vol}(B(0,R/2))}.$
One deduce by Plancherel
$\frac{\nu*\nu(B(0,R))}{\operatorname{Vol}(B(0,R))}\geq\frac{1}{2^{d}}\int_{V^{*}}h^{2}\left(\frac{x}{2R}\right)^{2}\hat{\nu}^{2}(x)\mathrm{d}x\geq\int_{B(0,A)}h^{2}\left(\frac{x}{2R}\right)^{2}\hat{\nu}^{2}(x)\mathrm{d}x,$
for $A>0$. Given such $A$, there is $R$ large enough such that the integrand is an increasing function of $R$ towards the function $\hat{\nu}^{2}$, and so by monotone convergence,
$\liminf_{R\rightarrow 0}\frac{\nu*\nu(B(0,R))}{\operatorname{Vol}(B(0,R))}\geq\int_{B(0,A)}\hat{\nu}^{2}(x)\mathrm{d}x.$
The conclusion follows by letting $A$ goes to +$\infty$+ ∞. ∎
{mdframed}
[backgroundcolor=white]
Lemma 3.10.
Let $E$ be a finite dimensional vector space. Let $\nu$ be a bounded measure on $V$, $\Sigma:V\rightarrow E$ be a measurable function and $f:E^{2}\rightarrow\mathbb{R}$ be a continuous function. Assume that
$\nu\left(\left\{x\in V\leavevmode\nobreak\ \middle|\leavevmode\nobreak\ f(\Sigma(x),\Sigma(x))>0\right\}\right)>0.$
Then there is $\alpha>0$ and a measurable subset $C$ of $V$ with positive measure such that
$\forall x,y\in C,\quad f(\Sigma(x),\Sigma(y))\geq\alpha.$
Proof.
Let $F:(x,y)\mapsto f(\Sigma(x),\Sigma(y))$. We define
$A=\left\{x\in V\leavevmode\nobreak\ \middle|\leavevmode\nobreak\ F(x,x)\geq 2\alpha\right\}.$
There are constants $\alpha,\varepsilon>0$ such that $\nu(A)>2\varepsilon$. By Lusin’s theorem, there is a compact subset $K\subset V$ such that $\nu(A\setminus K)\leq\varepsilon$ and such the restriction of $\Sigma$ to $A\cap K$ is uniformly continuous. Let $B=A\cap K$. Then $\nu(B)\geq\varepsilon$ and the restriction of $\Sigma$ to $B$ is a uniformly continuous function. The function $F$ is thus uniformly continuous on $B\times B$. We define $\omega$ its the modulus of continuity, meaning in particular that for $x,y\in B$ with $\|x-y\|\leq\eta$,
$F(x,y)\geq 2\alpha-\omega(\eta).$
Let us choose $\eta$ small enough so that $\omega(\eta)\geq\alpha$. The choice of a measurable subset $C\subset B$ of positive measure with diameter lower than $\eta$ concludes the proof. ∎
{mdframed}
[backgroundcolor=white]
Lemma 3.11.
let $\mu$ be a Hermitian measure on $V$ taking values in $\mathcal{S}q^{+}(W^{*})$, and $f$ be a symmetric element of $W^{*}\otimes W^{*}$. If $\operatorname{Im}(\mu)\not\subset C(f)$ then
$D^{-}\mu^{*2}(f)(0)>0.$
Proof.
Assume that $\mu$ has the representation $\mu=\Sigma\mathrm{d}\nu$, and define the function
$F(x,y)=\left(\Sigma(x)\otimes\Sigma(y)\right)(f).$
Since $\operatorname{Im}(\nu)\not\subset C(f)$, then thanks to the equivalence given by Lemma 3.3, one has
$\nu\left(\left\{x\in V\leavevmode\nobreak\ \middle|\leavevmode\nobreak\ F(x,x)>0\right\}\right)>0.$
By the previous Lemma 3.10, there is $\alpha>0$ and a subset $C$ of $V$ such that $\nu(C)>0$ and such that for $x,y\in C$, $F(x,y)\geq\alpha$. It implies that for $R>0$,
$\frac{\mu^{*2}(f)(B(0,R))}{\operatorname{Vol}(B(0,R))}$ $=\frac{1}{\operatorname{Vol}(B(0,R))}\int_{\|x-y\|\leq R}F(x,y)\mathrm{d}\nu(x)\mathrm{d}\nu(y)$
$\geq\frac{\alpha}{\operatorname{Vol}(B(0,R))}\int_{\|x-y\|\leq R}\mathds{1}_{C}(x)\mathds{1}_{C}(y)\mathrm{d}\nu(x)\mathrm{d}\nu(y)$
$\geq\alpha\frac{\nu_{C}*\nu_{C}(B(0,R))}{\operatorname{Vol}(B(0,R))},$
where $\nu_{C}$ is the measure on $V$ defined by the restriction of $\nu$ to $C$. Since $\nu(C)>0$, $\|\nu_{C}\|_{2}>0$ and the conclusion follows immediately from Lemma 3.9. ∎
{mdframed}
[backgroundcolor=white]
Lemma 3.12.
let $\mu$ be a Hermitian measure on $V$ taking values in $\mathcal{S}q^{+}(W^{*})$, and $f$ a Hermitian form on $W$ such that $\operatorname{Im}(\mu)\subset C(f)$. Assume that $\mu$ has the representation $\mu=\Sigma\mathrm{d}\nu$, where either
  • (1)
    $\nu$ is the Lebesgue measure of $V$ and $\Sigma\in L^{2}(V)$
  • (2)
    $\nu$ is the Lebesgue measure of a smooth compact hypersurface $M$ of $V$ and $\Sigma\in C^{1}(M)$.
Then
$D\mu^{*2}(f)(0)=0.$
Proof.
Given a Riemannian submanifold $M$ of $V$ of codimension $s$, let $\operatorname{dist}$ be its induced Riemannian distance, $\operatorname{Vol}_{M}$ is volume measure, and for $x\in M$ and $R>0$, let $B(x,M)$ the geodesic ball centered on $x$. If $\Sigma$ is a function defined on $M$, we define
$\Sigma_{R}:x\mapsto\frac{1}{\operatorname{Vol}_{M}(B_{M}(x,R))}\int_{B_{M}(x,R)}\Sigma(y)\mathrm{d}\nu(y).$
If $M=V$ and $\Sigma\in L^{2}(V)$, then for all $R>0$, one also has that $\Sigma_{R}\in L^{2}(V)$ and by a standard approximation technique, $\|\Sigma_{R}-\Sigma\|_{2}=o(1)$ as $R$ goes to 00. Similarly, if $M$ is a compact hypersurface and $\Sigma\in C^{2}(M)$, then using the Taylor expansion $\Sigma(y)=\Sigma(x)+\nabla_{x}\Sigma\cdot(y-x)+o(\|y-x\|)$, one has by symmetry
$\|\Sigma_{R}-\Sigma\|_{\infty}=o(R).$
In cases $(1)$ and $(2)$ we get the estimate for small $R$
$\|\Sigma_{R}-\Sigma\|_{2}=o(R^{s}).$
Note that for $s\geq 2$ the Taylor expansion of $\Sigma_{R}$ yields a non-zero second order term, which explained why we restricted ourselves to the case $s\geq 1$. Since $\operatorname{Im}(\mu)\subset C(f)$ then for almost all $x\in M$,
$\left(\Sigma(x)\otimes\Sigma(x)\right)(f)=0.$
There is a positive constant $C$ such that for $R$ small enough and $x,y\in M$ with $\operatorname{dist}(x,y)\leq M$,
$\frac{\operatorname{Vol}_{M}(B_{M}(x,R))}{\operatorname{Vol}(B(0,R/2))}\leq\frac{C}{R^{s}},\quad\text{and}\quad\operatorname{dist}(x,y)\leq 2\|x-y\|.$
One then has
$\frac{\mu^{*2}(f)(B(0,R/2))}{\operatorname{Vol}(B(0,R/2))}$ $=\frac{1}{\operatorname{Vol}(B(0,R/2))}\int_{\|x-y\|\leq\frac{R}{2}}\left(\Sigma(x)\otimes\Sigma(y)\right)(f)\mathrm{d}\nu(x)\mathrm{d}\nu(y)$
$\leq\frac{1}{\operatorname{Vol}(B(0,R))}\int_{\operatorname{dist}(x,y)\leq R}\left(\Sigma(x)\otimes\Sigma(y)\right)(f)\mathrm{d}\nu(x)\mathrm{d}\nu(y)$
$\leq\frac{C}{R^{d-s}}\left(\int_{M}\left(\Sigma(x)\otimes\Sigma_{R}(x)\right)(f)\mathrm{d}\nu(x)-\int_{M}\left(\Sigma(x)\otimes\Sigma(x)\right)(f)\mathrm{d}\nu(x)\right)$
$\leq\frac{C}{R^{s}}\|f\|^{2}\int_{M}\|\Sigma(x)\|\|\Sigma_{R}(x)-\Sigma(x)\|\mathrm{d}\nu(x)$
$\leq\frac{C}{R^{s}}\|f\|^{2}\|\Sigma\|_{L^{2}(\nu)}\|\Sigma-\Sigma_{R}\|_{L^{2}(\nu)}$
$\leq o(1),$
where the last line follows from Equation (3.4). The conclusion follows by passing to the limit as $R$ goes to 00. ∎
4 Kac–Rice formula, chaotic expansion and random waves
In this section, we first expose standard material concerning the Kac–Rice formula, in particular proving that one can formally embed its studies in light of the general framework of Section 1.1. We then compute the first terms of the chaotic expansion of the nodal volume. In a second part, we study the regularity of the convolution of spherical measures, which will be crucial for proving the fourth chaos asymptotics for the nodal volume of random waves.
4.1 The nodal measure of random fields
In this section we use the notation of the introduction. Let $(\delta_{u}^{\varepsilon})_{\varepsilon>0}$ be an approximation of the Dirac mass at 00 in $U$, denoted by $\delta$. We then define for $\phi$ a test function and $\lambda>0$ the approximated quantity
$Z_{\lambda}^{u,\varepsilon}(\phi)=\int_{U}\delta_{u}^{\varepsilon}(x)Z_{\lambda}^{x}(\phi)\mathrm{d}x.$
We define the gradient field $\nabla Y:V\rightarrow V\otimes U$. By the co-area formula, one has the a.s. equality
$Z_{\lambda}^{u,\varepsilon}(\phi)=\frac{1}{\lambda^{d/2}}\int_{V}\phi\left(\frac{v}{\lambda}\right)\delta_{u}^{\varepsilon}(Y(v))J(\nabla_{v}Y)\mathrm{d}v,$
where for an element $D\in U\otimes V$, $J(D)=\sqrt{\det(DD^{T})}$. {mdframed}[backgroundcolor=white]
Lemma 4.1.
We have the following convergence
$Z_{\lambda}^{u,\varepsilon}(\phi)\underset{\varepsilon\rightarrow 0}{\longrightarrow}Z_{\lambda}^{u}(\phi)$
a.s. and in $L^{2}(\Omega)$.
Proof.
The proof roughly follow the strategy in [Cratz-Leon]. One can show (see [Taylor-Adler] when $k=d$ or [Azais Wschebor]) that the function $Z_{\lambda}^{u}(\phi)$ is almost surely continuous in $u$. It implies the following a.s. convergence
$Z_{\lambda}^{u,\varepsilon}(\phi)\underset{\varepsilon\rightarrow 0}{\longrightarrow}Z_{\lambda}^{u}(\phi).$
To show that the convergence also holds in $L^{2}$, it suffices to show that $\mathbb{E}[Z_{\lambda}^{u,\varepsilon}(\phi)^{2}]$ converges to $\mathbb{E}[Z_{\lambda}^{u}(\phi)^{2}]$. From the result of [Gass-Stecconi], the nodal volume is a continuous functional of the covariance function of the field for the $L^{2}$ norm, and in particular the functions $u\rightarrow\mathbb{E}[Z_{\lambda}^{u}(\phi)^{2}]$ is continuous. From Fatou lemma, one has
$\mathbb{E}[Z_{\lambda}^{u}(\phi)^{2}]\leq\liminf_{\varepsilon\rightarrow 0}\mathbb{E}[Z_{\lambda}^{u,\varepsilon}(\phi)^{2}].$
Conversely, by Jensen one has
$\mathbb{E}[Z_{\lambda}^{u,\varepsilon}(\phi)^{2}]\leq\int_{\mathbb{R}^{k}}\mathbb{E}[Z_{\lambda}^{x}(\phi)^{2}]\delta^{\varepsilon}_{u}(x)\mathrm{d}x,$
and the right-hand side converges to $\mathbb{E}[Z_{\lambda}^{u}(\phi)^{2}]$ by continuity, when $\varepsilon$ goes to 00. ∎
Let the random field $X=(Y,\nabla Y)$, going from $V$ to the vector space
$W=U\times(U\otimes V)\simeq U\otimes(\mathbb{R}\times V),$
with associated covariance $\Omega$ and spectral measure $\mu$. Let $f:U\mapsto W$ be the tempered distribution, approximated by $f^{\varepsilon}$, defined for $x\in U$ and $D\in U\otimes V$ as
$f^{u}(x,D)=\delta_{u}(x)J(D)\quad\text{and}\quad f^{u,\varepsilon}(x,D)=\delta_{u}^{\varepsilon}(x)J(D)$
The rescaled nodal volume measure then has the following formal expression
$Z_{\lambda}(\phi)=\frac{1}{\lambda^{d/2}}\int_{V}\phi\left(\frac{v}{\lambda}\right)f(X(v))\mathrm{d}v,$
which must be understood as the limit in the $L^{2}$ norm given by Lemma 3.1. In particular, the random object $Z_{\lambda}(\phi)$ formally matches the general setting (1.1).
4.2 Chaotic expansion of the nodal measure
The $L^{2}$ convergence given by Lemma 4.1 implies that the function $f$ has a well-defined chaotic expansion of the form
$f:x\mapsto\sum_{q=0}^{+\infty}\frac{1}{q!}\langle f_{q},H_{x}^{q}\rangle,$
and one can apply the general formalism developed in Section 2.
Remark 4.2.
The zero set of $Y$ is independent of the choice of scalar product on $U$, and one can assume that $Y(0)$ is a standard Gaussian random variable. Similarly, a linear change of variables on $V$ modifies the nodal measure $dZ$ by a factor related to the determinant of the Jacobian transformation. One can assume that, for any fixed $u\in U$, $\nabla\langle Y(0),u\rangle$ is a standard Gaussian vector. Thus, the choice of scalar product on $V$ is arbitrary. However, one cannot assume that $X(0)=(Y(0),\nabla Y(0))$ is a standard Gaussian random variable. This is because the scalar product induced by $X(0)$ on $W=U\otimes(\mathbb{R}\times V)$ is generally different from the one inherited from any choice of scalar products on $U$ and $V$ due to a dimension argument. We must work with a rescaled version of $X$ and $f$ such that $X(0)$ becomes a standard Gaussian vector.
The semi-definiteness of $r$ implies that $r(-x)=r(x)^{T}$. Differentiating with respect to a vector $v$ at the point $x=0$, we find that $\operatorname{Cov}(Y(0),\nabla Y(0))$ is an anti-symmetric bilinear map. If $k=1$, this implies that $Y(0)\perp\nabla Y(0)$, and one can further assume that $X(0)$ is a standard Gaussian random variable, as explained in the previous paragraph. This fact partially explains why providing sufficient conditions to ensure the positivity of the variance is a more delicate task in the case $k>1$ and is difficult to treat in full generality. We show that, under natural additional constraints on the covariance function of the field, one can obtain asymptotic results. To compute the chaotic expansion explicitly, we use the following general lemma. {mdframed}[backgroundcolor=white]
Lemma 4.3.
Let $(X,Y)$ be a jointly Gaussian vector. Assume that $X$ is non-degenerate and let $\rho_{X}$ be its Gaussian density. Let $h$ be a continuous function with polynomial growth. Then
$\mathbb{E}[\delta(X)h(X,Y)]:=\lim_{\varepsilon\rightarrow 0}\mathbb{E}[\delta^{\varepsilon}(X)h(X,Y)]=\mathbb{E}[h(0,Y)|X=0]\rho_{X}(0).$
If $X$ and $Y$ are independent then
$\mathbb{E}[\delta(X)h(X,Y)]=\mathbb{E}[h(0,Y)]\rho_{X}(0).$
Proof.
The growth condition ensures that all the quantities are well defined. By definition of the conditional expectation
$\mathbb{E}[\delta^{\varepsilon}(X)h(X,Y)]$ $=\int\delta^{\varepsilon}(u)\mathbb{E}[h(x,Y)|X=x]\rho_{X}(x)\mathrm{d}x.$
Gaussian regression implies that the quantity $x\rightarrow\mathbb{E}[h(x,Y)|X=x]$ is well defined for all $x$ and continuous. The conclusion follows by letting $\varepsilon$ tend to zero. ∎
In light of Theorem 1.4 for the second chaos, we compute the spectral measure $\mu$ of the process $X=(Y,\nabla Y)$, and some quantities related to the second chaotic component $f_{2}$ of the tempered distribution $f$ defined in (4.1). {mdframed}[backgroundcolor=white]
Lemma 4.4.
$\mu=\psi\,\otimes\left[(1,2i\pi\xi)\otimes(1,-2i\pi\xi)\right].$
If $U=\mathbb{R}$ then
$\operatorname{Im}(\mu)=\left\{(1,2i\pi\xi)\leavevmode\nobreak\ \middle|\leavevmode\nobreak\ \xi\in\operatorname{supp}\psi\right\}\mathbb{C}.$
Proof.
The first point follows from the fact that the Fourier transform derivation into multiplication. In particular, for $q\geq 0$ the Fourier transform of the differential $D^{q}r$, taking values in $U\times V^{\otimes q}$, is given when defined by $\psi\otimes(2i\pi\xi)^{\otimes q}$. As for the second point, notice that the tensor $(1,2i\pi\xi)\otimes(1,-2i\pi\xi)$ is of rank one and its image is exactly the line directed by the vector $(1,2i\pi\xi)\in\mathbb{R}\times V$. The conclusion follows directly from the definition of the image of $\mu$, since $\mu\ll\psi$. ∎
We now compute quantities related to the second chaotic component $f_{2}$ of the formal function $f:(x,D)\mapsto\delta_{0}(x)J(D)$. Let $\rho(0)$ be the density at 00 of the vector $Y(0)$. One can assume, as explained in Remark 4.2, that $Y(0)$ is a standard Gaussian vector, and that for one fixed vector $v\in U$, the vector $Y_{v}=\nabla\langle Y(0),v\rangle$ is a standard Gaussian vector. We define the following quantities for a vector $w$ perpendicular to $v$, that are associated to the chaotic projection of order $0,1,2$ and $4$ of the volume of a $u$-level.
$q=0$ :
$\alpha=\mathbb{E}[f(X(0))]$
$q=1$ :
$\delta=\mathbb{E}[f(X(0))Y(0)],\quad\chi=\mathbb{E}[f(X(0))\nabla Y(0)]$
$q=2$ :
$\Lambda=\mathbb{E}[f(X(0))[Y(0)Y(0)^{T}-\mathrm{Id}]],\quad\Theta=\mathbb{E}[f(X(0))Y(0)\nabla Y(0)]$
$\beta_{u}=\frac{1}{d}\mathbb{E}[f(X(0))(\|Y_{v}\|^{2}-d)],\quad\kappa_{vw}=\mathbb{E}[f(X(0))Y_{v}^{T}Y_{w}],$
$q=4$ :
$\zeta_{v}=\frac{1}{d}\mathbb{E}[f(X(0))(\|Y_{v}\|^{2}-d)[Y(0)Y(0)^{T}-\mathrm{Id}]].$
{mdframed}
[backgroundcolor=white]
Lemma 4.5.
$\alpha\neq 0.$
$\delta=\alpha u,\quad\chi=0.$
$\quad\Theta=0,\quad\text{and}\quad\Lambda=\alpha(uu^{t}-\mathrm{Id}).$
If $Y_{v}$ is independent to $Y_{w}$, for $w\perp v$ then
$\kappa_{vw}=0,\quad\beta_{v}=\frac{\alpha}{d}\quad\text{and}\quad\zeta_{v}=-\frac{\alpha}{d}\mathrm{Id}.$
Proof.
The firsts statements are direct consequences of the symmetry of the Jacobian map, and Lemma 4.3. The second part of the proof borrows a nice trick from [Not21]. Notice first that
$\alpha=\rho(u)\mathbb{E}[J(\nabla Y(0))].$
Then the law of the vector $Y_{v}$ is invariant by rotation, which implies that
$\|Y_{v}\|\perp\frac{Y_{v}}{\|Y_{v}\|}.$
Moreover, the formula for the moments of a $\chi^{2}$ distribution implies that
$\frac{\mathbb{E}[\|Y_{v}\|^{3}]}{\mathbb{E}[\|Y_{v}\|]}=d+1.$
Now
$\beta_{v}$ $=\frac{1}{d}\rho(u)\mathbb{E}[J(Y_{v},(Y_{w})_{w})\|Y_{v}\|^{2}]-\alpha$
$=\frac{1}{d}\rho(u)\mathbb{E}\left[J\left(\frac{Y_{v}}{\|Y_{v}\|},(Y_{w})_{w}\right)\|Y_{w}\|^{3}\right]-\alpha$
$\overset{\perp}{=}\frac{1}{d}\rho(u)\mathbb{E}[J(Y_{v},(Y_{w})_{w})]\frac{\mathbb{E}[\|Y_{v}\|^{3}]}{\mathbb{E}[\|Y_{v}\|]}-\alpha$
$=\frac{d+1}{d}\alpha-\alpha$
$=\frac{\alpha}{d}.$
At last, by independence of $Y$ and $\nabla Y$,
$\zeta_{v}=\frac{\beta_{v}\Lambda}{\alpha}=\frac{\alpha}{d}(uu^{t}-\mathrm{Id}).$
4.3 Self-convolution of hypersurface measures
The goal of this subsection is to derive formulas for the convolution of continuous measures supported on the sphere. These computations could certainly be approached from a probabilistic perspective, specifically by considering the distribution of a uniform random walk, for which precise results concerning the probability of returning close to zero after a small number of steps exist; see, for instance, [Bor12]. We define the convolution symbol $\;\bar{*}\;$ as
$\mu\;\bar{*}\;\nu=\mu\;\bar{*}\;(\nu(-\,\cdot)).$
If the measure $\nu$ is symmetric w.r.t. the origin, it coincides with the usual convolution operation. We say that a function $f$ defined on the ball $B(0,R)$ in $\mathbb{R}^{d}$ is polar-continuous if the function $\tilde{f}:(r,\theta)\mapsto f(r\theta)$ is continuous on $[0,R]\times\mathcal{S}^{d-1}$. The typical example of a polar continuous (but not continuous) function is the mapping $x\mapsto x/\|x\|$. {mdframed}[backgroundcolor=white]
Lemma 4.6.
Let $\sigma$ be the uniform probability measure on the sphere of radius $1$. Then $\sigma*\sigma$ has a density w.r.t the Lebesgue measure given by
$\sigma*\sigma(x)=C_{d}\frac{1}{\|x\|}(4-\|x\|^{2})^{\frac{d-3}{2}}\mathds{1}_{B(0,2)}(x),$
for some normalization constant $C_{d}$. If $\rho_{1},\rho_{2}$ are continuous functions on the sphere, then the measure $(\rho_{1}\sigma)\;\bar{*}\;(\rho_{2}\sigma)$ has a polar-continuous density w.r.t the measure $\sigma*\sigma$. Precisely, if $\mu_{\theta}$ designs the probability measure of the $d-2$ dimensional sphere of radius $1$ and normal vector $\theta$, then
$\lim_{(r,\theta)\rightarrow(0,\theta_{0})}\frac{\mathrm{d}(\rho_{1}\sigma)\;\bar{*}\;(\rho_{2}\sigma)}{\mathrm{d}\sigma*\sigma}(r,\theta)=\int\rho_{1}(y)\rho_{2}(y)\mathrm{d}\mu_{\theta_{0}}(y),$
Proof.
Let $M$ and $N$ two compact hypersurface of $\mathbb{R}^{d}$ with respective Lebesgue measure $\sigma_{M}$ and $\sigma_{N}$. Let $\rho_{M}$ and $\rho_{N}$ be two smooth functions on $M$ and $N$, respectively. We assume that $M$ and $N$ are transverse, meaning that for all $x\in M\cap N$, the unit normal vectors of the two hypersurfaces at point $x$ form an angle $\theta_{x}\in]0,\pi[$. By compactness, there is $\eta>0$ such that for all $u\in B(0,\eta)$, $M+u$ is also transverse to $N$. Moreover, $M\cap N$ is a submanifold of dimension $d-2$, and by considering a $\varepsilon$-neighborhood of $M$ and $N$ respectively, one can observe, taking the limit as $\varepsilon$ goes to 00, that $(\rho_{M}\sigma_{M})\;\bar{*}\;(\rho_{N}\sigma_{N})$ has a density in $B(0,\eta)$ with respect to the Lebesgue measure on $\mathbb{R}^{d}$ such that
$(\rho_{M}\sigma_{M})\;\bar{*}\;(\rho_{N}\sigma_{N})(0)=\int_{M\cap N}\frac{\rho_{M}(x)\rho_{N}(x)}{|\sin(\theta_{x})|}\mathrm{d}\sigma_{N\cap M}(x).$
A bit of geometry shows that two unit spheres whose centers are shifted by a vector $x$ with norm $r\in]0,2[$ intersects everywhere at an angle $\theta_{r}$ such that
$\sin\left(\frac{\theta_{r}}{2}\right)=\frac{r}{2},\quad\text{so that}\quad\sin(\theta_{r})=2r\sqrt{1-\left(\frac{r}{2}\right)^{2}}.$
The intersection of the two spheres is a $d-2$ dimensional sphere $B_{x}$ of radius $\sqrt{1-\left(\frac{r}{2}\right)^{2}}$. Gathering this fact and Formula (4.3), we deduce that the convolution $\sigma*\sigma$ has a density w.r.t. Lebesgue measure given by
$\sigma*\sigma(x)=C_{d}\frac{1}{\|x\|}(4-\|x\|^{2})^{\frac{d-3}{2}}\mathds{1}_{B(0,2)}(x),$
and more generally
$(\rho_{1}\sigma)\;\bar{*}\;(\rho_{2}\sigma)(x)=\left(\int_{B_{x}}\rho_{1}(y)\rho_{2}(y)\mathrm{d}\mu_{x}(y)\right)\sigma*\sigma(x),$
where $\mu_{x}$ is the uniform probability measure on the $d-2$ dimensional sphere $B_{x}$. Passing in polar coordinates $x=(r,\theta)$, the sphere $B_{x}$ converges, as $r$ goes to 00 and $\theta$ goes to $\theta_{0}$, to the $d-2$ sphere $B_{\theta_{0}}$ of radius $1$ and normal vector $\theta_{0}$, and
$\lim_{(r,\theta)\rightarrow(0,\theta_{0})}\int_{B_{x}}\rho_{1}(y)\rho_{2}(y)\mathrm{d}\mu_{x}(y)=\int_{B_{\theta_{0}}}\rho_{1}(y)\rho_{2}(y)\mathrm{d}\mu_{\theta_{0}}(y).$
Remark 4.7.
The key assumption regarding the sphere is that it is everywhere positively curved. The curvature implicitly appears as the differential of the normal vector at the start of the proof. A similar degeneracy in the function $\sigma*\sigma$ could also be observed for more general positively curved manifolds $M$, near the point $x=0$ and along the boundary of the set of points $x$ such that $M\cap(M+x)\neq\emptyset$. By linking the regularity of a function with the decay of its Fourier transform, this provides an alternative heuristic explanation —beyond the usual stationary phase approach— of why the Fourier transform of a set with positively bounded curvature exhibits favorable decay properties (see the definition of test functions in Section 1.1).
{mdframed}
[backgroundcolor=white]
Lemma 4.8.
Let $\sigma$ be the uniform probability measure on the sphere of radius $1$. Then the measure $\sigma^{*4}$ has a density w.r.t the Lebesgue measure such that for some positive constant $C_{d}$,
$\sigma^{*4}(x)\simeq\left\{\begin{array}[]{llll}C_{2}|\log(\|x\|)|\;\text{ if }d=2,\\ C_{d}\;\qquad\quad\quad\;\text{ if }d\geq 3,\\ \end{array}\right.$
In dimension $2$, let $\omega_{1}$ and $\omega_{2}$ be polar continuous functions on the closed ball $B(0,2)$. Then there are positive constants $C$, $C^{\prime}$ such that
$\lim_{x\rightarrow 0}\frac{1}{|\log(\|x\|)|}(\omega_{1}\sigma^{*2})\;\bar{*}\;(\omega_{2}\sigma^{*2})(x)=C\int_{\mathcal{S}^{1}}\omega_{1}(0,\theta)\omega_{2}(0,\theta)\mathrm{d}\theta+C^{\prime}\int_{\|y\|=2}\omega_{1}(y)\omega_{2}(y)\mathrm{d}y.$
Proof.
The previous Lemma 4.6 implies for $d\geq 3$ that the measure $\sigma*\sigma$ is square integrable. In particular, the measure $\sigma^{*4}$ has a continuous density, and $\sigma^{*4}(0)=\|\sigma*\sigma\|_{2}^{2}>0$. When $d=2$, the function $\sigma*\sigma$ is not square-integrable. One has
$(\omega_{1}\sigma)\;\bar{*}\;(\omega_{2}\sigma)(x)=C\int_{\mathbb{R}^{d}}\frac{\mathds{1}_{B(0,2)}(x+y)\mathds{1}_{B(0,2)}(y)\omega_{1}(x+y)\omega_{2}(y)}{\|y\|\|x+y\|\sqrt{4-\|y\|^{2}}\sqrt{4-\|x+y\|^{2}}}\mathrm{d}y.$
We must understand the singular loci : near the point $y=0$, and near the sphere $\|y\|=2$. Let $x=(\varepsilon,\theta)$. Near $y=0$ we make the substitution $y=\varepsilon u$. The singularity comes from the term $1/\|u\|\|\theta+u\|$ that integrates as a $|\log(\varepsilon)|$ on $B(0,2/\varepsilon)$. One can then isolate the mass of the singularity for instance on the domain
$D_{\varepsilon}=B\left(0,\frac{2}{\varepsilon|\log(\varepsilon)|}\right)\setminus B(0,|\log(\varepsilon)|).$
On this domain one has $\theta+u\simeq u$ and $\varepsilon u\rightarrow 0$, so that as $\varepsilon$ goes to 00,
$(\rho_{1}\sigma)\;\bar{*}\;(\rho_{2}\sigma)(x)\simeq C\int_{D_{\varepsilon}}\frac{\omega_{1}(\varepsilon u)\omega_{2}(\varepsilon u)}{4\|u\|^{2}}\mathrm{d}y\simeq C|\log(\varepsilon)|\int_{\mathcal{S}^{1}}\omega_{1}(0,\theta)\omega_{2}(0,\theta)\mathrm{d}\theta.$
By a similar argument one can obtain the same logarithmic scale near the sphere $\|y\|=2$. ∎
5 Proofs of the main theorems
In this section, we prove the main theorem stated in introduction. The first section is devoted to the proof of the general theorems, which are relatively straightforward consequences of the lemmas stated in Section 3. The second subsection exposes the proof related to the nodal measure, which follow from the material developed in Section 4. In the following, we will use the notation developed in the previous subsection.
5.1 Proofs of the general theorems
In this section, we prove the general theorems stated in introduction, namely 1.1, Corollaries 1.2 and 1.3 and Theorem 1.4. For $u,v\in V$, one has from Equation (7)
$\operatorname{Var}(f(X(u)),f(X(v)))=\sum_{q=1}^{+\infty}\mathbb{E}[\pi_{q}(f)(X(u))\pi_{q}(f)(X(v))]=\sum_{q=1}^{+\infty}\frac{1}{q!}(\Omega(u-v))^{\otimes q}(f_{q},f_{q}).$
One deduce that
$\operatorname{Var}(Z_{\lambda}(\phi))$ $=\sum_{q=1}^{+\infty}\frac{1}{\lambda^{d}}\int_{V^{2}}\phi\left(\frac{u}{\lambda}\right)\phi\left(\frac{v}{\lambda}\right)\mathbb{E}[\pi_{q}(f)(X(u))\pi_{q}(f)(X(v))]\mathrm{d}u\mathrm{d}v$
$=\sum_{q=1}^{+\infty}\operatorname{Var}(Z_{\lambda}^{(q)}(\phi)),$
proving the first part of the lemma. As for the second part, the stationarity implies the relation
$\operatorname{Var}(Z_{\lambda}^{(q)}(\phi))=\frac{1}{q!}\int_{V}(\Omega(u-v))^{\otimes q}(f_{q},f_{q})\left(\frac{1}{\lambda^{d}}\phi\left(\frac{\,.\,}{\lambda}\right)*\phi\left(\frac{-\,.\,}{\lambda}\right)\right)(v)\mathrm{d}v$
Using Plancherel theorem, one obtain
$\operatorname{Var}(Z_{\lambda}^{(q)}(\phi))=\frac{1}{q!}\int_{V^{*}}\lambda^{d}|\hat{\phi}(\lambda\xi)|^{2}\mathrm{d}\widehat{\Omega^{\otimes q}(f_{q},f_{q})}(\xi)=\frac{1}{q!}\int_{V^{*}}\gamma_{\lambda}(\xi)\,\mathrm{d}\mu^{*q}(f_{q})(\xi).$
The proof of Corollary 1.2 is a direct application of Lemma 3.8 applied to the measure $\mu^{*q}(f_{q})$. As for the next Corollary 1.3, the integrability assumption on the density $\Sigma$ implies that $\Sigma^{*q}$ is a bounded and continuous function. So that
$\lim_{\lambda\rightarrow+\infty}\operatorname{Var}(Z_{\lambda}^{(q)}(\phi))=\Sigma^{*q}(f_{q})(0).$
The conclusion follows from the convolution formula
$\Sigma^{*q}(0)=\int_{\xi_{1}+\ldots+\xi_{q}=0}\bigotimes_{k=1}^{q}\Sigma(\xi_{k})\mathrm{d}\xi.$
When $q=1$ and $q=2$, these expressions reduces to the formulas stated.
5.2 Proofs of the variance of the nodal measure
The first section is devoted to the proof of Theorem 1.5. The second subsection is devoted to the proof Theorem 1.7 and Theorem 1.8. The last subsection is devoted to the proof of Theorem 1.9.
5.2.1 Proof of Theorem 1.5
The proof of the asymptotic normality exactly follows the same strategy as [Nua20], and is based on the Fourth moment theorem [Nua05]. The argument for the cancellation of the contractions is completely identical. The only difference is that in our setting, $f$ is a tempered distribution and one cannot use directly the usual Arcones inequality [Arc94] to upper bound the variance of $Z_{\lambda}(\phi)$, and we detail here the proof for the variance bound. We use the notations of Section 2.3. By assumption on the field, one has $\Omega(v)\in E^{*}(\eta)$, where $\eta$ is the Gaussian measure on $W$. We define, for $v\neq 0$ the quantity
$\mathbb{E}[f(0)f(v)]=(f\otimes f)(\rho_{\Omega}).$
It follows from [Gas24] applied with $p=2$, or directly from the standard divided difference trick presented in the introduction of that paper, that there is a constant $C$ such that for $v\in B(0,1)$
$\mathbb{E}[f(0)f(v)]\leq C.$
By Kac–Rice formula
$\operatorname{Var}(Z_{\lambda}(\phi))$ $=\int_{V\setminus\{0\}}\frac{1}{\lambda^{d}}\phi*\phi\left(\frac{v}{\lambda}\right)\mathbb{E}\left[f(X(0))f(X(v))\right]\mathrm{d}v+\mathbb{E}[Z_{\lambda}(\phi)]^{2}\mathds{1}_{k=d}$
Since $\Omega$ converges to 00 as $\|v\|\rightarrow+\infty$, by compactness one can find a compact $A$ of $E^{*}(\eta)$ such that for all $v\in V\setminus B(0,1)$, one has $\Omega(v)\in A$. Then by Equation (2.3), there is a constant $C_{q}$ such that for $v\in V\setminus B(0,1)$,
$\left|\mathbb{E}\left[f(X(0))f(X(v))\right]-\sum_{k=0}^{q-1}\mathbb{E}\left[f_{k}(X(0))f_{k}(X(v))\right]\right|\leq C_{q}\|\Omega\|^{q}.$
The function $\frac{1}{\lambda^{d}}\phi*\phi(\cdot/\lambda)$ uniformly converges to the constant function $1$, and the convergence of the variance of $Z_{\lambda}^{(q+)}(\phi)$ follows by dominated convergence, gathering Equation (5.2.1) and (5.2.1).
5.2.2 Proof of Theorem 1.7 and Theorem 1.8
Let $\Sigma=\widehat{\Omega}$ be the spectral density of the random process $X=(Y,\nabla Y)$ that takes values in the space $W=U\times(U\otimes V)\simeq U\otimes(\mathbb{R}\times V)$. It follows from Corollary 1.7 that the limit expression for the variance is given by the quantity
$\lim_{\lambda\rightarrow+\infty}\operatorname{Var}(Z_{\lambda}^{(2)}(\phi))=\frac{1}{2}\int_{V^{*}}\operatorname{Tr}(f_{2}\Sigma(\xi)f_{2}\Sigma(\xi))\mathrm{d}\xi.$
According to Lemma 4.4, the spectral density $\Sigma$ has expression
$\Sigma(\xi)=\psi(\xi)\,\otimes\left[(1,2i\pi\xi)\otimes(1,-2i\pi\xi)\right].$
$\bullet$ Assume the hypothesis $(H1)$:
It follows from Lemma 4.5 that the second chaotic component $f_{2}$ has the expression, in an orthogonal basis on $W$
$f_{2}=\mathrm{Id}\otimes\begin{pmatrix}[c|c]-\alpha&0\\ \hline\cr 0&\frac{\alpha}{d}\,\mathrm{Id}\end{pmatrix}.$
Using the definition of the trace, or equivalently the mixed product property, and the second part of Lemma 3.2 applied to the rank one matrix $(1,2i\pi\xi)\otimes(1,-2i\pi\xi)$, one gets
$\operatorname{Tr}\left(f_{2}\Sigma(\xi)f_{2}\Sigma(\xi)\right)$ $=\operatorname{Tr}(\sigma(\xi)^{2})\left((1,2i\pi\xi)^{T}\begin{pmatrix}[c|c]-\alpha&0\\ \hline\cr 0&\frac{\alpha}{d}\,\mathrm{Id}\end{pmatrix}(1,-2i\pi\xi)\right)^{2}$
$=\alpha^{2}\left(-1+\frac{4\pi^{2}}{d}\,\|\xi\|^{2}\right)^{2}\|\psi(\xi)\|_{2}^{2}.$ (17)
Since the function $\psi$ is in $L^{2}$ and is non identically zero, it cannot be supported on the zero set of a non-zero quadratic equation. Therefore, this last quantity is not the zero function and
$\lim_{\lambda\rightarrow+\infty}\operatorname{Var}(Z_{\lambda}^{(2)}(\phi))>0.$
Note that the proof didn’t make use of Theorem 1.4 because the computation are explicit. If we remove the hypothesis of square integrability on $\mu$, then let $\mathrm{d}\psi=\varsigma(\xi)\mathrm{d}\nu$ be a representation of $\sigma$. We deduce the following representation for the spectral measure $\mu$
$\mathrm{d}\mu=\Theta(\xi)\mathrm{d}\nu,\quad\text{with}\quad\Theta(\xi)=\varsigma(\xi)\,\otimes\left[(1,2i\pi\xi)\otimes(1,-2i\pi\xi)\right]\mathrm{d}\nu.$
In the case $k=1$ the isotropic cone of $f_{2}$ and the image of $\mu$ are given explicitly from Lemma 4.4 by
$C(f_{2})=\left\{(1,2i\pi\xi)\leavevmode\nobreak\ \middle|\leavevmode\nobreak\ \|\xi\|=d\right\}\mathbb{C}\quad\text{and}\quad\operatorname{Im}(\mu)=\left\{(1,2i\pi\xi)\leavevmode\nobreak\ \middle|\leavevmode\nobreak\ \xi\in\operatorname{supp}\sigma\right\}\mathbb{C},$
and $\operatorname{Im}(\mu)\subset C(f_{2})$ if and only if the measure $\sigma$ is supported on the sphere of radius $\sqrt{d}/2\pi$, i.e. if $Y$ is a random wave. The conclusion of Theorem 1.8 directly follows from the two points of Theorem 1.4, applied to the sphere as a smooth compact hypersurface of $V^{*}$. In the case $k>1$, the computation of the isotropic cone and the image are not convenient and we follow a more direct approach. The same computation as in Equation (17) shows that
$\operatorname{Tr}\left(f_{2}\varsigma(\xi)f_{2}\varsigma(\xi)\right)=\alpha^{2}\left(-1+\frac{4\pi^{2}}{d}\,\|\xi\|^{2}\right)^{2}\|\varsigma(\xi)\|_{2}^{2}.$
If $Y$ is not a random wave then this quantity is positive on a set of positive $\nu$-measure, and the conclusion follows from the proof Lemma 3.11. Conversely, if $Y$ is a regular random wave then this quantity cancels for all $\xi$ in the support of $\nu$ and the conclusion of Theorem 1.8 follows from the proof of Lemma 3.12, and in particular the second point.
$\bullet$ Assume the hypothesis $(H2)$:
If $U=\mathbb{R}$, the result follows directly from the previous hypothesis. If $k>1$ one can assume that, up to a change of coordinates on $V$, $\nabla\langle Y,u\rangle$ is a standard Gaussian. Let $\psi_{u}$ be the spectral density of $\langle Y,u\rangle$. The spectral density $\Sigma$ is block diagonal with block $\Sigma_{u}$ and $\Sigma_{u^{\perp}}$, where
$\Sigma_{u}(\xi)=\psi_{u}(\xi)\left[(1,2i\pi\xi)\otimes(1,-2i\pi\xi)\right].$
By independence, it follows from Lemma 4.5 that the second chaotic component is block diagonal with blocks $f_{2,u}$ and $f_{2,u^{\perp}}$, where
$f_{2,u}=\begin{pmatrix}[c|c]-\alpha&0\\ \hline\cr 0&\alpha M\end{pmatrix}$
and $M$ is a matrix (it is actually a diagonal matrix with trace $d$). As in the proof for hypothesis $(H1)$, one has
$\operatorname{Tr}\left(f_{2}\Sigma(\xi)f_{2}\Sigma(\xi)\right)\geq\left(-\alpha+4\pi^{2}\frac{\alpha}{d}\,\xi^{T}M\xi\right)^{2}|\psi_{u}(\xi)|^{2},$
and the conclusion follows in a similar fashion as in the previous case.
$\bullet$ Assume the hypothesis $(H3)$:
Let $F$ a continuous Gaussian field from $V$ to $\mathbb{R}$ with covariance function $q$ and spectral measure $\omega$, such that $Y=\nabla F$. The field $Y$ is taking values in $V$ and one can assume by a proper rescaling that $r(0)=\operatorname{Hess}q(0)=\mathrm{Id}$. Then, refining expression (5.2.2), one has
$\Sigma(\xi)=\omega(\xi)\left[(2i\pi\xi)\otimes(-2i\pi\xi)\right]\otimes\left[(1,2i\pi\xi)\otimes(1,-2i\pi\xi)\right].$
As for the second chaotic component $f_{2}$, it can be written in an orthonormal basis as
$f_{2}=\begin{pmatrix}[c|c]-\alpha\,\mathrm{Id}&0\\ \hline\cr 0&H\end{pmatrix},$
for some matrix $M$. As pointed out in Remark 4.2 we have to work with the rescaled field $\tilde{X}=\Pi X$. Because $Y(0)$ is already a standard Gaussian vector, independent from $\nabla Y(0)$, the mapping $\Pi$ can be chosen so that it only acts on the part $V\otimes V$ of $W$. The mapping $\tilde{\Sigma}$ has rank one, and it follows again from Lemma 3.2 that
$\operatorname{Tr}\left(f_{2}\tilde{\Sigma}(\xi)f_{2}\tilde{\Sigma}(\xi)\right)=\omega(\xi)^{2}\left(-4\pi^{2}\xi^{2}\alpha+(\Pi(2i\pi\xi)\otimes(2i\pi\xi))^{T}H\Pi(-2i\pi\xi)\otimes(-2i\pi\xi)\right)^{2}.$
The conclusion follows in a similar fashion as in the previous case.
$\bullet$ Conclusion:
The Hermite rank of the nodal volume functional is two, and $r\in L^{2}(V)$. It follows from Theorem 1.5 applied with $q=2$ that the limiting variance of $Z_{\lambda}$ exists, is finite and under any of the three hypotheses of Lemma 1.7, is positive thanks to the lower bound
$\operatorname{Var}(Z_{\lambda}(\phi))\geq\operatorname{Var}(Z_{\lambda}^{(2)}(\phi))>0.$
The CLT for $Z_{\lambda}(\phi)$ again follows from Theorem 1.5.
Remark 5.1.
For $u$-level with $u\neq 0$, one can prove by similar arguments that there is no second chaos cancellation for the isotropic random waves model. Indeed, Equation (5.2.2) becomes, thanks to Lemma 4.5,
$\operatorname{Tr}\left(f_{2}\Sigma(\xi)f_{2}\Sigma(\xi)\right)=\alpha^{2}\left(\|u\|^{2}-1+\frac{4\pi^{2}}{d}\|\xi\|^{2}\right)^{2}+(k-1)\alpha^{2}\left(-1+\frac{4\pi^{2}}{d}\|\xi\|^{2}\right)^{2}>0.$
Since the measure $\mu$ is not square integrable, the limiting variance explodes and one can explicitly compute the asymptotics, as it has been done in [Ros16], and prove the CLT by standard methods.
5.2.3 Proof of Theorem 1.9
Let $\mu$ be the spectral measure of the process $(Y,\nabla Y)$ taking values in the space of Hermitian forms on $W=U\otimes(\mathbb{R}\times V)$. We want to apply Theorem 1.4 with $q=2$, i.e. to show that $\operatorname{Im}(\mu^{*2})\notin C(f_{4})$ and conclude that the variance converges to a limit constant. The case $d=2$ is a bit different because of the logarithmic singularity in Lemma 4.8, but the proof follows the same line as in the case $d\geq 3$ with minor adaptations that will be detailed.
From Lemma 4.4, the spectral measure $\mu$ has a density w.r.t the spherical measure $\sigma$ given by
$\mathrm{d}\mu(\xi)=\mathrm{Id}\otimes\left[(1,2i\pi\xi)\otimes(1,-2i\pi\xi)\right]\mathrm{d}\sigma(\xi),$
Thanks to Lemma 4.6, the measure $\mu*\mu$ has a density $\Sigma$ w.r.t the measure $\sigma*\sigma$, which is bounded and polar-continuous. In light of 1.4 with $q=2$, we want to show that $\operatorname{Im}(\Sigma)\notin C(f_{4})$. Since $\Sigma$ is polar-continuous, it is enough to show the following lemma, whose proof is postponed to the end of the present proof. {mdframed}[backgroundcolor=white]
Lemma 5.2.
For $\theta\in\mathcal{S}^{d-1}$, one has
$\operatorname{Im}(\Sigma(0,\theta))\not\subset C(f_{4}).$
We distinguish the cases $d=2$ and $d\geq 3$.
Case $d\geq 3$:
In that case, $\Omega\in L^{4}$ and it follows from Theorem 1.5 that the variance of $Z_{\lambda}^{(4+)}(\phi)$ converges to a finite constant, which is positive thanks to Lemma 5.2 and Lemma 1.4, and satisfies a CLT. Since the variance of $Z_{\lambda}^{(2)}(\phi)$ converges to zero, then $Z_{\lambda}^{(2)}(\phi)$ converges in probability to zero and the CLT for $Z_{\lambda}(\phi)$ follows from Slutsky’s Theorem.
Case $d=2$:
The variance of the fourth chaos has the expression
$\operatorname{Var}(Z_{\lambda}^{(4)}(\phi))$ $=\frac{1}{24}\int_{\mathbb{R}^{d}}\left(|\hat{\phi}|^{2}\right)_{\lambda}(\xi)\,\mathrm{d}(\mu*\mu)^{*2}(f_{4})(\xi)$
$=\frac{1}{24}\int_{\mathbb{R}^{d}}\left(|\hat{\phi}|^{2}\right)_{\lambda}(\xi)\,\mathrm{d}\left[\Sigma(\sigma^{*2})*\Sigma(\sigma^{*2})\right](f_{4})(\xi)$
The variance asymptotics is explicit thanks to Lemma 4.8. By integration of equivalences, it is given for some positive constant $C,C^{\prime}$ by
$\lim_{\lambda\rightarrow+\infty}\frac{\operatorname{Var}(Z_{\lambda}^{(4)}(\phi))}{\log(\lambda)}=C\int_{\mathcal{S}^{1}}\Sigma(0,\theta)^{\otimes 2}(f_{4})\mathrm{d}\theta+C^{\prime}\int_{\|\xi\|=\frac{\sqrt{d}}{\pi}}\Sigma(\xi)^{\otimes 2}(f_{4})\mathrm{d}x.$
Note that both terms are non-negative. From Lemma 5.2 and Lemma 3.3, the integrand in the first term is positive and the conclusion follows. Note that one can completely explicit the Hermitian form $\Sigma$ (see for instance the following proof of Lemma 5.2), and the fourth chaotic projection $f_{4}$ (by similar computation as in Lemma 4.5) to explicit the limiting constant. The CLT for $Z_{\lambda}^{(4)}(\phi)$ follows directly from the fourth Moment theorem, since $\Omega\in L^{6}$ implies that the contractions tends to zero. The variance of $Z_{\lambda}^{(6+)}(\phi)$ converges to a constant, and so does the variance of $Z_{\lambda}(\phi)-Z_{\lambda}^{(4)}(\phi)$. The CLT for $Z_{\lambda}(\phi)$ follows from Slutsky’s Theorem.
Proof of Lemma 5.2 :
Proof.
Assume first that $U=\mathbb{R}$, i.e. $k=1$. The mapping $\Sigma$ acts on the space of Hermitian forms on the complexification of the space $W=\mathbb{R}\times V$. Let $e_{0},e_{1},\ldots,e_{d}$ be a basis of ×$\mathbb{R}\times V$.The choice of $\theta$ is unimportant by isotropy of the problem and one can assume that $\theta=e_{2}$. We write the mapping $\Sigma(0,e_{2})$ in the induced basis on $V$. For $0\leq i\leq j\leq d$ and $0\leq k\leq l\leq d$, we define
$\Sigma_{(ij)(kl)}=\Sigma(0,e_{2})(\operatorname{Sym}(e_{i}\otimes e_{j}),\operatorname{Sym}(e_{k}\otimes e_{l})),$
All these quantities can be computed thanks to the second part of Lemma 4.6 : let $\rho_{0}=1$ and for $1\leq j\leq$, let $\rho_{j}:x\mapsto 2i\pi\langle x,e_{j}\rangle$. If we define $B_{1}$ the $d-2$-sphere of radius $1$, normal vector $e_{2}$ and uniform probability measure $\sigma_{2}$, then up to a normalization constant $C$ one has
$\Sigma_{(ij)(kl)}=C\int_{B_{1}}\rho_{i}(x)\rho_{j}(x)\overline{\rho_{k}(x)}\overline{\rho_{l}(x)}\mathrm{d}\;\sigma_{2}(x).$
One can explicitly compute all the coefficient of $\Sigma$ but it won’t be necessary : all we need is to observe that $\Sigma_{(01)(01)}>0$, and for $(ij)\neq(01)$, one has $\Sigma_{(01)(ij)}=0$ by symmetry. It implies that the unit vector $\operatorname{Sym}(e_{0}\otimes e_{1})$ is in the image of $\Sigma(0,\theta)$. Then
$(f_{4})(\operatorname{Sym}(e_{0}\otimes e_{1}),\operatorname{Sym}(e_{0}\otimes e_{1}))$ $=f_{4}(e_{0}\otimes e_{0}\otimes e_{1}\otimes e_{1})$
$=\mathbb{E}[f(X(0))(Y(0)^{2}-1)((\partial_{1}Y(0))^{2}-1)]$
$=\zeta$
where the last equality follows from Lemma 4.5. Since $\zeta\neq 0$, it implies that $\operatorname{Sym}(e_{0}\otimes e_{1})\notin C(f_{4})$ and thus that $\operatorname{Im}(\Sigma(0,\theta))\not\subset C(f_{4})$, which proves the lemma in the case $k=1$.
For general $U$, the mapping $\Sigma$ acts on the space of Hermitian forms on the complexification of the space $W=U\otimes(\mathbb{R}\times V)$. Let $u_{1},\ldots,u_{k}$ be a basis of $U$. Notice first that the action of mapping $\Sigma$ can be decomposed as the sum of its actions on each space $W_{i}=\{u_{i}\}\otimes(\mathbb{R}\times V)$, by independence of the coordinates. It then suffices to apply the reasoning above reasoning for $k=1$, to the restriction of $\Sigma$ to any of these spaces to show that the unity vector $\operatorname{Sym}((u_{i}\otimes e_{0})\otimes(u_{i}\otimes e_{1}))$ is in the image of $\Sigma(0,e_{2})$, but not in the isotropic cone $C(f_{4})$, and the conclusion follows. ∎
[…]